### v2

# Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation

Abstract.
Existing solutions for bundle recommendation(BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. However, bundle-item(B-I) affiliation will vary dynamically in real scenarios. For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments. Our empirical study demonstrates that the performance of mainstream BR models will fluctuate or even decline regarding item-level variability. This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation(RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by BR model to represent bundle theme via a forward-reverse process. In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings. In particular, the residual connection in our residual approximator significantly enhances item-level bundle embeddings generation ability of BR models. Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%. Codes and datasets are available at https://anonymous.4open.science/r/RDiffBR.
Bundle Recommendation, Diffusion Model, Dynamic Variability Scenario
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; New York,USAisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Recommender systems
1. INTRODUCTION
With the rapid development of online platforms, bundle recommendation(BR) has become an important tool for enhancing user experience and promoting sales. Specifically, users are provided with one-stop services by adding a group of related items to a bundle (such as booklists, playlists, fashion sets, and meal sets in takeout services) while improve sales efficiency and revenue.
Bundle recommendation usually learns from the user-bundle (U-B) interactions, user-item (U-I) interactions and bundle-item (B-I) affiliations to recommend a list of bundles to users. However, in real scenarios B-I affiliations may undergo dynamic changes due to factors such as seasonal variations, shifts in user preferences, or inventory adjustments, while the theme of the bundle remains. As illustrated in Figure 1, Bundle 1 initially includes ‘T-shirt’, ‘pants’, ‘shoes’ and ‘watch’ with the theme of ‘casual outfit’.
Figure 1. An example shows bundle-item affiliation is changed while bundle theme is relatively stable.
[…]
In practice, platforms can add the ‘hat’ to align with summer promotions or remove the ‘watch’ based on user feedback. An intuitive idea is to regard it as a new bundle. However, such item-level changes will not cause bundle theme shifting although bundle theme is jointly determined by the items within the bundle. As shown in Figure 1, the theme of Bundle 2 and Bundle 3 remain ‘casual outfit’ by item variation. It is advisable to give priority to reusing the identification information of the original bundle because this dynamic variability maintains the theme of bundle.
In recent years, lots of works have focused on bundle recommendation and achieved good results. Including factorization models represented by LIRE (Liu et al., 2014), EFM (Cao et al., 2017), attention-based models represented by DAM (Chen et al., 2019b), Attlist (He et al., 2019) and GNN-based models represented by BGCN (Chang et al., 2020), CrossCBR (Ma et al., 2022). However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline in B-I dynamic variability scenarios we have mentioned above. For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases(see Section 3.2 for details). This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations regarding item-level.
In this paper, we make the first attempt to address the issue of item-level dynamic variability in bundle recommendation and propose a novel Residual Diffusion for Bundle Recommendation(RDiffBR) model-agnostic generative framework to assist the model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by BR model to represent bundle theme. In the forward process, RDiffBR injects controlled Gaussian noises into item-level bundle embeddings. In the reverse process, RDiffBR iteratively removes noises by a designed residual approximator. RDiffBR jointly trains this residual approximator and BR model via above forward-reverse process. In the inference stage, to generate the effective item-level bundle embeddings, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios rather than using pure noise as the starting point and conduct bundle recommendation. Experiments on six bundle models and four public datasets from different domains show that RDiffBR significantly improves the performance of bundle models in dynamic B-I variability scenario.
The main contributions of this work are summarized as follows:
$\bullet$ We propose a generative residual diffusion framework for bundle recommendation to address item-level dynamic variability issue.
$\bullet$ Our residual approximator in diffusion can significantly enhance item-level bundle embeddings generation ability of BR models.
$\bullet$ Extensive experiments on four datasets from different domains show that RDiffBR enhances the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
Figure 2. Overview of the proposed RDiffBR framework.
2. Methodology
Generally speaking, there is an approache to deal with the item-level dynamic variability. Utilize side information (such as item attributes) to introduce additional features but requires architectural modifications to the original model. In contrast, our RDiffBR does not require additional information. The following will give our technical details.
2.1. Problem Definition
$U(|U|=M)$ is denoted as the user set, $B(|B|=N)$as the bundle set, and $I(|I|=O)$ as the item set. The user-bundle interactions, user-item interactions, and bundle-item affiliations are represented as $X_{M\times N}=\{(u,b)\mid u\in U,b\in B\}$, $Y_{M\times O}=\{(u,i)\mid u\in U,i\in I\}$ and $Z_{N\times O}=\{(b,i)\mid b\in B,i\in I\}$ with a binary value at each entry respectively. The value 1 indicates an observed interaction of user-item or user-bundle, or affiliation of bundle-item. The item-level dynamic variability problem can be formulated as follows:
Input: The user-bundle interactions $X_{M\times N}$, user-item interactions $Y_{M\times O}$, and bundle-item affiliations $Z_{N\times O}$.
Output: A well-trained bundle recommendation estimates the probabilty $\hat{y}_{ub}$ of a user $u$ interacting with a bundle $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$.
Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation. And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
2.2. RDiffBR Framework
Considering the successful experience of diffusion as a generative model and the sparsity of input data, our RDiffBR framework introduces a residual diffusion model in the latent space of interactions. As shown in Figure 2, after BR model utilizes an encoder to get bundle-level user/bundle embeddings $E^{BL}_{u},E^{BL}_{b}$ and item-level embeddings user/bundle embeddings $E^{IL}_{u},E^{IL}_{b}$ from $\{X,Y,Z\}$. Since $E^{IL}_{b}$ represents the bundle theme generated by the BR model, we iteratively process $E^{IL}_{b}$ through a forward-backward process, jointly optimizing the BR model and the residual approximator module that we have designed based on the task. In the inference stage, RDiffBR applies this optimized forward-reverse process to $E^{IL}_{b}$ from dynamic variability B-I affiliations $Z^{*}$ to obtain new $E^{IL}_{b}$ for bundle recommendation, enhancing BR model’s ability for this scenario.
2.2.1. Residual Diffusion.
$\bullet$ Forward process. Given item-level bundle embeddings $E^{IL}_{b}$ from a BR model’s encoding we can obtain a bundle’s item-level bundle embedding $e^{IL}_{b}$ and set $e_{0}=e^{IL}_{b}$. As shown in the bottom of residual diffusion module in Figure 2, Gaussian noises are continuously incorporated into $e_{0}$ with adjustable scales and steps:
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}).$
Based on the reparameterization (Ho et al., 2020) and the additivity of independent Gaussian noises (Luo, 2022), we can directly obtain $e_{t}$ from $e_{0}$.
(2) $q(e_{t}|e_{0})=\mathcal{N}(e_{t},\sqrt{\bar{\alpha}_{t}}e_{0},(1-\bar{\alpha_{t}})\boldsymbol{I}),$
where $\bar{\alpha_{t}}=\prod_{t=1}^{t}\alpha_{t},\alpha_{t}=1-\beta_{t},\beta_{t}\in(0,1)$ controls the Gaussian noise scales added at each step $t$. To regulate the added noises in $e_{1:T}$, we follow  (Wang et al., 2023) using a linear variance noise schedule in the forward process:
(3) $1-\bar{\alpha}_{t}=s\cdot\left[\alpha_{\min}+\frac{t-1}{T-1}(\alpha_{\max}-\alpha_{\min})\right],$
where T is the total forward step and s $\in$ (0, 1) controls the noise scale. The condition $\alpha_{min}$ $<$ $\alpha_{max}$in (0, 1) indicates the upper and lower limits of the added noise.
$\bullet$ Reverse process. After obtaining the noisy embedding $e_{T}$ by forward process, we iteratively denoise this embedding in reverse processing as shown in the top of residual diffusion module in Figure 2. In each reverse step, a residual approximator module to denoise them.
(4) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\Sigma_{\theta}(\hat{e}_{t},t)),$
where $\hat{e}_{t}$ is denoised embedding in the reverse step t, $\theta$ and are the learnable parameters of residual approximator. This module are executed iteratively in the reverse process until the reconstruction of final embedding $\hat{e}_{0}$
$\bullet$ Residual approximator. We design a residual approximator module based on our task to denoise the noisy embedding in each reverse step as shown in the right of residual diffusion module. The diffusion training aims to approximate the distribution $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})$ with the tractable posterior distribution $q(e_{t-1}|e_{t},e_{0})$ in the reverse process. Through Bayes’ theorem, we can derive:
(5) $q(e_{t-1}|e_{t},e_{0})\propto\mathcal{N}(e_{t-1};\tilde{\mu}(e_{t},e_{0},t),\sigma^{2}(t)\boldsymbol{I}),\mathrm{~{}where}$
$\begin{cases}\tilde{\mu}(e_{t},e_{0},t)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}e_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}e_{0}\\ \sigma^{2}(t)=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},&\end{cases}$
$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (Zhao et al., 2024) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
(6) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~{}where}$
$\mu_{\theta}(\hat{e}_{t},t)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\hat{e}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\tilde{e}_{0},$
$\tilde{e}_{0}$ is the predicted $e_{0}$ since the distribution of $e_{0}$ is unknown in the reverse process. We employ the multi-layer perceptron (MLP) to reconstruct $e_{0}$ in the residual approximator. Specifically, for embedding $e_{t}$ in the reverse step t:
(7) $\tilde{e}_{0}=\delta f_{\theta}(\hat{e}_{t},t)+(1-\delta)e_{t^{\prime}},$
where $\tilde{e}_{0}$ is the predicted value of $e_{0}$ by residual approximator with parameter $\theta$. $\delta$ is residual connection weight. $e_{t^{\prime}}$ is noisy embedding in the forward step $t^{\prime}$. Step information t is encoded by sinusoidal positional encoding (Ho et al., 2020), and these two inputs are concated as input.
2.2.2. Loss Function.
To optimize the forward-reverse process, it is necessary to minimize the variational lower bound of the predicted embedding. According to the KL divergence based on the multivariate Gaussian distribution, the reconstruction loss of within one training iteration can be expressed as:
(8) $\mathcal{L}_{r}$ $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{e}_{t-1}|\mathbf{e}_{t},\mathbf{e}_{0}\right)\|p_{\theta}\left(\mathbf{e}_{t-1}|\mathbf{e}_{t}\right)\right)\right]$
$=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[\frac{1}{2\sigma^{2}(t)}\left[\|\mu_{\theta}(\mathbf{e}_{t},t)-\mu(\mathbf{e}_{t},\mathbf{e}_{0},t)\|_{2}^{2}\right]\right].$
Furthermore, by substituting Eq.(5) and Eq.(6) into Eq.(8) and simplifying (Wang et al., 2023) we can drive:
(9) $\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{t\sim\mathcal{U}(1,T)}\mathbb{E}_{q}\left[||e_{0}-\hat{e}_{\theta}(e_{t},e_{t^{\prime}},t)||_{2}^{2}\right],$
where $\mathcal{U}$ means uniformly sampling t from $\{1,2,...,T\}$. The final loss function of RDiffBR consists of: a bundle model loss and a reconstruction loss for reverse process. We have designed a loss balancing factor $\lambda$ to adjust the weight of reconstruction loss and jointly train the bundle model and residual approximator together:
(10) $\mathcal{L}=\mathcal{L}_{BR\mathrm{~{}model}}+\lambda\mathcal{L}_{\mathrm{diff}}.$
The training step of residual approximator is illustrated in Algorithm 1. Specifically, Lines 2-5 represent the forward process, Line 6 describes the reverse process using residual approximator, and Lines 7-8 address optimization.
Input: interaction data $X$, item-level bundle embedding $e_{0}$, diffusion step $T$, randomly initialized residual approximator MLP $f_{\theta}$.
1 repeat
2       Sample a batch of U-B interactions $\bar{X}\subset X$.
3
4      for $all\ (u,b)\in\bar{X}$  do
5             Sample $t\sim\mathcal{U}(1,T)$;
6
7            Compute $e_{t}$ given $e_{0}$ and $t$ via $q(e_{t}|e_{0})$ in Eq.(2);
8
9            Reconstruct $\tilde{e}_{0}$ through $f_{\theta}$ in Eq.(7);
10            Calculate $\mathcal{L}$ by Eq.(10);
11
12            Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
13       end for
14
15until converged;
Output: optimized $\theta.$
Algorithm 1 Residual approximator training process
2.2.3. Inference.
In the inference stage, we take the item-level bundle embedding $e^{*}_{0}$ in the B-I dynamic variability scenario as input. Subsequently, RDiffBR continuously introduces noise into $e^{*}_{0}$ in the sequence $e^{*}_{0}\to e^{*}_{1}\to...\to e^{*}_{T^{\prime}}$ for $T^{\prime}$ steps in the forward process and then we set $\hat{e}_{{T^{\prime}}}=e^{*}_{T^{\prime}}$ to execute reverse process $\hat{e_{T^{\prime}}}\to\hat{e}_{{T^{\prime}}-1}\to...\to\hat{e}_{0}$ to generate new item-level bundle embedding $\hat{e}_{0}$. The reverse ignores the variance and uses $\hat{e}_{t-1}=\mu_{\theta}(\hat{e}_{t},e^{*}_{0},t)$ for directly inference following (Liang et al., 2018). We replace the original $e^{*}_{0}$ with $\hat{e}_{0}$, and then conduct bundle recommendation following the original way of the bundle model. Finally enhancing the ranking of bundles in dynamic variation scenarios. The inference procedure of residual approximator is illustrated in Algorithm 2. Specifically, Line 1 represent the forward process, Lines 2-5 represent the reverse process using a residual approximator to obtain ideal item-level bundle embedding.
1
Input: item-level bundle embedding $e_{0}^{*}$ from B-I dynamic variability scenario, sampling step $T^{{}^{\prime}}$, well-trained residual approximator MLP $f_{\theta}$.
2
3Compute $e_{T^{{}^{\prime}}}^{*}$ given $e_{0}^{*}$ and $T^{{}^{\prime}}$ via $q(e_{T^{{}^{\prime}}}^{*}|e_{0}^{*})$ in Eq.(2).
4
5for $t=T^{{}^{\prime}},...,1$ do
6       Reconstruct $\tilde{e}_{0}$ through $f_{\theta}$ in Eq.(7);
7      Compute $\hat{e}_{t-1}$ from $\hat{e}_{t}$ and $\tilde{e}_{0}$ in Eq.(6);
8 end for
9
Output: item-level bundle embedding $e_{0}$.
Algorithm 2 Inference process
3. EXPERIMENTS
Our experiments aim to answer the following research questions:
$\bullet$ RQ1: How does RDiffBR perform on different datasets of backbone bundle models under various degrees of B-I variability?
$\bullet$ RQ2: What is the contributions of residual connection in RDiffBR to its overall performance?
$\bullet$ RQ3: How does the efficiency of RDiffBR?
3.1. Experimental Settings
$\bullet$ Datasets. Four publicly bundle datasets from different domains are summarized in Table 1. Youshu (Chen et al., 2019b) for book list recommendation; NetEase (Cao et al., 2017) for music playlist recommendation; iFashion (Chen et al., 2019a) for fashion outfit recommendation; In addition, we introduce another meal recommendation dataset $\mathrm{MealRec+}$ (Li et al., 2024a), where the meal consisted of individual course items is treated as bundle.
Table 1. Dataset Statistics.
Dataset #U #I #B #U-I #U-B #B-I #Avg.I/B #Dens.B-I
Youshu 8,039 32,770 4,771 138,515 51,377 176,667 37.03 0.11%
NetEase 18,528 123,628 22,864 1,128,065 302,303 1,778,838 77.8 0.06%
iFashion 53,897 42,563 27,694 2,290,645 1,679,708 106,916 3.86 0.01%
$\mathrm{MealRec+}$ 1,575 7,280 3,817 151,148 46,767 11,451 3 0.77%
$\bullet$ Baselines. We implement RDiffBR on six representative BR models with official code released by their authors. BGCN (Chang et al., 2020) utilizes graph convolutional networks to separately learn the representations of bundle view and item view. HyperMBR (Ke et al., 2023a) models two views in hyperbolic space to learn their accurate representations. MIDGN (Zhao et al., 2022) refines the user-item/bundle-item graph of different intents from multiple views. CrossCBR (Ma et al., 2022) improves two separate views by constrastive learning. MultiCBR (Ma et al., 2024a) first fuses the multi-view representations into a unified one, and then adopts a contrastive loss on the unified representations. BundleGT (Wei et al., 2023) learns the strategy-aware user and bundle representation in bundle recommendation. There are still some earlier research outcomes (for instance, BundleNet (Deng et al., 2020)) that have been studied to be less effective than the aforementioned methods. Therefore, we do not take them into consideration.
$\bullet$ Experimental process. According to previous works (Li et al., 2024a) (Ma et al., 2022), the training, validation and testing sets of user-bundle interactions $X$ are randomly splited with specific ratio. It is not feasible to simply add new bundles in bundle-item affiliations $Z$ under the B-I variability assumptions. Therefore, we adjust existing bundle datasets to match this scenario: Let $\{{\mathcal{E}_{1},\mathcal{E}_{2},…,\mathcal{E}_{10}}\}$ denote a uniform partition of $Z$ into ten disjoint subsets. Bundle models are trained and validated on $Z_{\mathrm{train\&val}}=\cup_{i=1}^{5}\mathcal{E}_{i}$. In testing phase, we design $Z_{\mathrm{test}}$ by adjusting $Z_{\mathrm{train\&val}}$ with adding or removing subsets:
(11) $Z_{\mathrm{test}}=\begin{cases}Z_{\mathrm{train\&val}}\cup\left(\mathcal{\cup}_{i=6}^{5+\rho}\mathcal{E}_{i}\right)&\mathrm{if}~{}\rho>0\\ Z_{\mathrm{{train\&val}}}&\mathrm{if~{}\rho=0}\\ Z_{\mathrm{train\&val}}\setminus\left(\mathcal{\cup}_{i=6-|\rho|)}^{5}\mathcal{E}_{i}\right)&\mathrm{if~{}\rho<0}\end{cases}$
Where $\cup$ indicates set union and $\setminus$ indicates set difference. $\rho\in\{-4,-3,-2,-1,0,1,2,3,4,5\}$ reflects level of B-I variation. $\rho>0$ indicates adding set $\Delta^{+}$ to $Z_{\mathrm{test}}$, $\rho<0$ indicates removing set $\Delta^{-}$ from $Z_{\mathrm{test}}$ and $\rho=0$ indicates $Z_{\mathrm{test}}$ invariability. The well-trained bundle models predict user-bundle interactions $X$ on $Z_{test}$. The processing of the B-I affiliation is shown in Figure 3. In this way, we can control the degree of B-I affiliaitons variation and simulate the scenario of B-I dynamic variability.
Figure 3. The processing of the B-I affiliation to match the scenario of B-I dynamic variability.
$\bullet$ Hyperparameter Settings We set the hyperparameters of six backbone models from their authors and optimize RDiffBR with the Adam optimizer with a learning rate of 0.001. The time steps $T$ and sampling steps $T^{\prime}$ are set to values ranging from [1,200]. The loss balance factor $\lambda$ is chosen from (0,5), while the residual connection weight $\delta$ is selected from (0,1). The number of MLP layers is selected within 1, 2, 4, 8. The hidden size of the MLP layers are chosen from [8,4096]. Details of other hyperparameters can be found in our released code. All the models are trained using two GTX 3090.
3.2. Overall Performance (RQ1)
Table 2. Performance comparison between backbone models and our RDiffBR on four datasets($\rho$=-4, -2, 0, 2, 4), while $\rho<0$ indicates $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\setminus\Delta^{-}$, $\rho>0$ indicates adding set $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\cup\Delta^{+}$, and $\rho=0$ indicates $Z_{\mathbf{train\&val}}$ invariability. Improv. indicates the relative improvement of a backbone model+RDiffBR over the itself.
Dataset Youshu NetEase
$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
BGCN 0.1897 0.1047 0.2254 0.1278 0.2367 0.1380 0.2214 0.1247 0.2138 0.1204 0.0581 0.0309 0.0634 0.0333 0.0643 0.0339 0.0646 0.0341 0.0622 0.0332
+RDiffBR 0.2178 0.1278 0.2391 0.1415 0.2453 0.1466 0.2470 0.1423 0.2407 0.1396 0.0614 0.0321 0.0657 0.0340 0.0662 0.0343 0.0668 0.0345 0.0649 0.0340
improv. 14.83% 22.15% 6.06% 10.74% 3.61% 6.25% 11.57% 14.11% 12.60% 16.03% 5.56% 3.98% 3.63% 1.99% 2.89% 1.14% 3.47% 1.07% 4.42% 2.54%
MIDGN 0.2388 0.1373 0.2557 0.1472 0.2598 0.1496 0.2578 0.1495 0.2566 0.1489 0.0540 0.0278 0.0625 0.0330 0.0636 0.0337 0.0639 0.0337 0.0640 0.0337
+RDiffBR 0.2498 0.1441 0.2621 0.1498 0.2645 0.1515 0.2652 0.1517 0.2643 0.1516 0.0568 0.0293 0.0649 0.0340 0.0661 0.0347 0.0667 0.0348 0.0672 0.0350
improv. 4.62% 4.98% 2.53% 1.72% 1.79% 1.32% 2.87% 1.49% 3.00% 1.82% 5.21% 5.27% 3.89% 2.91% 3.96% 2.92% 4.45% 3.25% 4.95% 4.14%
HyperMBR 0.1464 0.0953 0.2126 0.1290 0.2649 0.1581 0.2167 0.1285 0.1897 0.1162 0.0588 0.0313 0.0675 0.0354 0.0717 0.0377 0.0689 0.0363 0.0661 0.0345
+RDiffBR 0.1789 0.1109 0.2229 0.1327 0.2669 0.1594 0.2280 0.1347 0.2128 0.1275 0.0627 0.0332 0.0694 0.0369 0.0739 0.0386 0.0704 0.0369 0.0667 0.0350
improv. 22.20% 16.42% 4.85% 2.87% 0.75% 0.83% 5.20% 4.87% 12.14% 9.75% 6.67% 5.97% 2.88% 4.41% 3.07% 2.51% 2.25% 1.79% 0.85% 1.50%
CrossCBR 0.2677 0.1584 0.2763 0.1626 0.2776 0.1640 0.2788 0.1649 0.2800 0.1656 0.0715 0.0386 0.0776 0.0419 0.0794 0.0431 0.0802 0.0436 0.0798 0.0434
+RDiffBR 0.2744 0.1640 0.2814 0.1659 0.2827 0.1663 0.2816 0.1664 0.2816 0.1662 0.0772 0.0419 0.0818 0.0436 0.0826 0.0443 0.0829 0.0444 0.0826 0.0445
improv. 2.52% 3.57% 1.84% 2.02% 1.85% 1.42% 1.01% 0.92% 0.56% 0.40% 7.99% 8.50% 5.36% 4.04% 4.09% 2.67% 3.43% 1.70% 3.50% 2.51%
MultiCBR 0.2770 0.1649 0.2769 0.1649 0.2791 0.1652 0.2767 0.1641 0.2770 0.1648 0.0843 0.0463 0.0867 0.0472 0.0876 0.0475 0.0880 0.0479 0.0879 0.0479
+RDiffBR 0.2831 0.1674 0.2840 0.1677 0.2824 0.1673 0.2825 0.1669 0.2842 0.1675 0.0859 0.0469 0.0879 0.0478 0.0881 0.0480 0.0883 0.0482 0.0893 0.0487
improv. 2.20% 1.55% 2.56% 1.71% 1.19% 1.25% 2.10% 1.71% 2.60% 1.61% 1.91% 1.32% 1.37% 1.27% 0.57% 1.18% 0.40% 0.75% 1.60% 1.52%
BundleGT 0.2687 0.1590 0.2872 0.1683 0.2877 0.1694 0.2873 0.1697 0.2837 0.1679 0.0729 0.0367 0.0869 0.0465 0.0874 0.0468 0.0874 0.0468 0.0871 0.0465
+RDiffBR 0.2829 0.1653 0.2927 0.1708 0.2922 0.1707 0.2939 0.1712 0.2903 0.1702 0.0759 0.0384 0.0900 0.0479 0.0896 0.0479 0.0897 0.0477 0.0895 0.0476
improv. 5.27% 3.96% 1.90% 1.49% 1.57% 0.79% 2.30% 0.93% 2.31% 1.40% 4.06% 4.75% 3.49% 3.16% 2.52% 2.42% 2.67% 1.86% 2.76% 2.37%
Dataset iFashion $\mathbf{{MealRec+}}$
$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
BGCN 0.0473 0.0331 0.0560 0.0395 0.0602 0.0426 0.0607 0.0428 0.0604 0.0427 0.1068 0.1888 0.1213 0.2094 0.1325 0.2247 0.1326 0.2250 0.1375 0.2323
+RDiffBR 0.0488 0.0341 0.0568 0.0398 0.0606 0.0428 0.0608 0.0429 0.0607 0.0429 0.1180 0.2078 0.1293 0.2180 0.1408 0.2401 0.1387 0.2371 0.1388 0.2384
improv. 3.09% 2.91% 1.55% 0.58% 0.61% 0.45% 0.16% 0.23% 0.57% 0.47% 10.07% 10.55% 4.08% 6.53% 6.82% 6.26% 5.38% 4.57% 2.63% 0.92%
MIDGN 0.0372 0.0257 0.0549 0.0380 0.0672 0.0482 0.0428 0.0279 0.0441 0.0294 0.0820 0.1564 0.1017 0.1816 0.1121 0.1908 0.1129 0.1919 0.1115 0.1880
+RDiffBR 0.0426 0.0290 0.0628 0.0437 0.0763 0.0553 0.0534 0.0348 0.0545 0.0363 0.0826 0.1604 0.1075 0.1930 0.1179 0.2050 0.1145 0.2012 0.1146 0.2027
improv. 14.55% 12.83% 14.36% 14.87% 13.53% 14.65% 24.66% 24.83% 23.48% 23.77% 2.60% 0.82% 6.29% 5.69% 7.44% 5.13% 4.87% 1.48% 7.82% 2.79%
HyperMBR 0.0448 0.0336 0.0687 0.0494 0.0815 0.0580 0.0773 0.0545 0.0744 0.0521 0.0859 0.1498 0.0939 0.1595 0.1069 0.1778 0.1057 0.1781 0.1106 0.1887
+RDiffBR 0.0462 0.0349 0.0692 0.0504 0.0830 0.0595 0.0797 0.0567 0.0773 0.0550 0.0874 0.1585 0.0995 0.1760 0.1122 0.1944 0.1205 0.1988 0.1247 0.2144
improv. 3.21% 4.09% 0.81% 2.01% 1.75% 2.69% 3.01% 4.01% 3.92% 5.42% 5.82% 1.70% 10.33% 5.90% 9.34% 4.94% 11.61% 14.00% 13.62% 12.82%
CrossCBR 0.0624 0.0491 0.0877 0.0683 0.0999 0.0772 0.1046 0.0799 0.1081 0.0831 0.0612 0.1048 0.0819 0.1417 0.0938 0.1688 0.1084 0.2041 0.1504 0.2599
+RDiffBR 0.0720 0.0554 0.0928 0.0716 0.1027 0.0794 0.1060 0.0812 0.1088 0.0834 0.0681 0.1203 0.0926 0.1531 0.1026 0.1788 0.1241 0.2243 0.1611 0.2669
improv. 15.36% 12.87% 5.75% 4.78% 2.83% 2.82% 1.31% 1.64% 0.60% 0.32% 14.80% 11.40% 8.01% 13.15% 5.95% 9.37% 9.89% 14.48% 2.69% 7.12%
MultiCBR 0.0672 0.0588 0.1080 0.0905 0.1297 0.1064 0.1343 0.1101 0.1381 0.1133 0.0989 0.1782 0.1147 0.1841 0.1305 0.2092 0.1476 0.2376 0.1625 0.2592
+RDiffBR 0.0694 0.0608 0.1094 0.0916 0.1313 0.1080 0.1350 0.1111 0.1385 0.1139 0.1066 0.1889 0.1217 0.1945 0.1375 0.2164 0.1524 0.2436 0.1633 0.2598
improv. 3.29% 3.31% 1.23% 1.16% 1.20% 1.47% 0.53% 0.97% 0.27% 0.58% 6.04% 7.83% 5.63% 6.10% 3.44% 5.40% 2.51% 3.19% 0.25% 0.50%
BundleGT 0.0444 0.0352 0.0690 0.0517 0.0954 0.0724 0.0899 0.0653 0.0960 0.0707 0.0732 0.1281 0.0923 0.1556 0.1093 0.1761 0.1217 0.2134 0.1579 0.2585
+RDiffBR 0.0456 0.0361 0.0701 0.0530 0.0970 0.0740 0.0909 0.0660 0.0970 0.0712 0.0863 0.1512 0.1082 0.1770 0.1315 0.2077 0.1414 0.2346 0.1707 0.2733
improv. 2.68% 2.58% 1.64% 2.34% 1.61% 2.10% 1.02% 1.04% 1.10% 0.71% 18.03% 17.88% 13.75% 17.20% 17.97% 20.35% 9.91% 16.21% 5.73% 8.05%
We present the results of Recall@20 and NDCG@20 for five levels of B-I variations in Table 2. We show the results of Recall@40 and NDCG@40 on two datasets for ten levels of B-I variations in Figure 4. The following observations can be made:
$\bullet$ The perfomance of BR models when meeting item-level dynamic variability. When $\rho<0$ ($\mathbf{Z_{test}}$ removes $\Delta^{-}$), the performance of bundle models is generally lower than that under $\rho=0$ in most cases. The greater the reduction in B-I, the more significant the performance degradation. When $\rho>0$ ($\mathbf{Z_{test}}$ adds $\Delta^{+}$), some models exhibit performance decline, others maintain comparable performance, while a few show improvement. This indicates that BR models possess capability to handle B-I variations to some extent, yet they tend to fail under drastic reductions of B-I. Notably, HyperMBR performs the worst, where both B-I adding and removing frequently lead to performance deterioration. This phenomenon may stem from hyperbolic space embeddings being more sensitive to distance errors compared to Euclidean space. In contrast, models like MIDGN, CrossCBR, and MultiCBR exhibit relatively stable performance, potentially attributed to their adoption of contrastive learning that leverages multi-view information through cooperative learning. Additionally, BundleGT demonstrates enhanced adaptability for dynamic B-I variations by modeling B-I affiliation via attention mechanisms.
$\bullet$ The performance of RDiffBR dealing with item-level dynamic variability. Our experiments reveal that RDiffBR consistently enhances the performance of six backbone BR models across four datasets from different domains under different B-I variation levels. When $\rho<0$ ($\mathbf{Z_{test}}$ removes $\Delta^{-}$), The performance of BR models+RDiffBR outperform BR models with the improvement up to 22.15% at most. When $\rho>0$ ($\mathbf{Z_{test}}$ adds $\Delta^{+}$), RDiffBR also enhances the performance of BR models, reaching up to 24.83% at most. In some cases, RDiffBR can even enable BR models to outperform its performance under $\rho=0$, which highlights the powerful generative capability of RDiffBR. Furthermore, performance improvements is observed even under $\rho=0$, which demonstrates RDiffBR has certain effect even in the B-I remain scenario. Overall, RDiffBR can enhance the performance of BR models with different architectures and metric spaces under different variation scenarios. This remarkable performance improvement can be attributed to RDiffBR’s generation capability, which reconstructs item-level bundle embeddings in B-I change scenarios to better align with bundle themes rather than overfitting to the B-I affiliations observed during training.
Figure 4. Performance comparisons of the backbone models and our RDiffBR under different levels of B-I variations in terms of Recall@40 and NDCG@40.
3.3. Ablation Study (RQ2)
To assess whether residual connections contribute to improving model performance, we set a variant RDiffBR w/o R that omits the residual connection in Eq.(7) and conduct a comprehensive ablation study on three BR models across three datasets, with the results shown in Table 3.
Table 3. Ablation study where bold indicates the best method, and underlined indicates the second best method($\rho=-3,3$).
Method Youshu NetEase $\mathrm{MealRec+}$
R@20 N@20 R@20 N@20 R@20 N@20
BGCN $\underline{0.2214}$ $\underline{0.1280}$ 0.0621 $\underline{0.0328}$ 0.1962 0.1122
BGCN+RDiffBR 0.2366 0.1398 0.0647 0.0336 0.2159 $\underline{0.1240}$
BGCN+RDiffBR w/o R 0.2158 0.1151 $\underline{0.0624}$ 0.0323 $\underline{0.2118}$ 0.1245
CrossCBR $\underline{0.2742}$ 0.1623 0.0767 0.0413 0.1229 0.0720
CrossCBR+RDiffBR 0.2789 0.1646 0.0810 0.0433 $\underline{0.1329}$ $\underline{0.0798}$
CrossCBR+RDiffBR w/o R 0.2740 $\underline{0.1628}$ $\underline{0.0782}$ $\underline{0.0418}$ 0.1508 0.0825
BundleGT $\underline{0.2836}$ $\underline{0.1656}$ $\underline{0.0838}$ $\underline{0.0438}$ 0.1398 0.0841
BundleGT+RDiffBR 0.2908 0.1693 0.0864 0.0450 $\underline{0.1616}$ $\underline{0.0982}$
BundleGT+RDiffBR w/o R 0.2519 0.1454 0.0520 0.0267 0.2313 0.1416
Method Youshu NetEase $\mathrm{MealRec+}$
R@20 N@20 R@20 N@20 R@20 N@20
BGCN 0.2129 0.1197 $\underline{0.0632}$ $\underline{0.0336}$ $\underline{0.2298}$ $\underline{0.1350}$
BGCN+RDiffBR 0.2394 0.1374 0.0660 0.0343 0.2390 0.1404
BGCN+RDiffBR w/o R $\underline{0.2337}$ $\underline{0.1229}$ 0.0627 0.0321 0.2089 0.1221
CrossCBR $\underline{0.2795}$ $\underline{0.1653}$ $\underline{0.0801}$ $\underline{0.0435}$ $\underline{0.2395}$ $\underline{0.1295}$
CrossCBR+RDiffBR 0.2816 0.1664 0.0827 0.0445 0.2474 0.1440
CrossCBR+RDiffBR w/o R 0.2763 0.1639 0.0798 0.0426 0.1843 0.1052
BundleGT $\underline{0.2858}$ $\underline{0.1691}$ $\underline{0.0871}$ $\underline{0.0466}$ $\underline{0.2394}$ 0.1416
BundleGT+RDiffBR 0.2925 0.1711 0.0895 0.0477 0.2589 0.1593
BundleGT+RDiffBR w/o R 0.2519 0.1454 0.0520 0.0267 0.2312 $\underline{0.1417}$
Figure 5. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=-3$.
Figure 6. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=3$.
From this table, we can draw the following conclusions: In most cases, the performance of using RDiffBR is superior to that of not using it or using RDiffBR w/o R. Using RDiffBR w/o R may lead to worse performance than RDiffBR, and in some cases it may even be worse than only using backbone BR model. Figure 6-6 show the visualization results of CrossCBR, CrossCBR+RDiffBR and CrossCBR+RDiffBR w/o R under two B-I varation levels for item-level bundle embeddings $E^{IL}_{B}$ in Youshu dataset. It can be observed that $E^{IL}_{B}$ generated by CrossCBR after B-I change is more dispersed than original $E^{IL}_{B}$ and more compact $E^{IL}_{B}$ distribution which retains the basic shape can be obtained after using CrossCBR+RDiffBR. This indicates that $E^{IL}_{B}$ of the BR model cannot effectively represent the bundle theme distribution due to the B-I change and RDiffBR enables the model to obtain more bundle information at the item-level rather than overfitting to the existing b-I affiliations. While using CrossCBR+RDiffBR w/o R, $E^{IL}_{B}$ distribution changes little so the effect is limited, which is consistent with the experimental results obtained in Table 3. We also conducted a case study to explore the specific situations of RDiffBR and RDiffBR w/o R(see detail in section3.5). Overall, RDiffBR outperforms backbone models, highlighting the effectiveness of RDiffBR. Residual connection plays an important role in RDiffBR, and merely using a diffusion model cannot generate better item-level bundle embeddings.
Furthermore, our task can be approached without considering the Out-Of-Distribution(OOD) which means the distribution of the training data and the test data is inconsistent. Although different B-I affiliations are used during model testing, the bundle theme remains, which means items within a bundle remain to follow the similar distribution before and after the change. This can also be verified by Figure 6-6 that the shapes of two $E^{IL}_{B}$ distributions under different B-I variation $\rho$ are similar.
3.4. Impact of Hyperparameters
Figure 7. The impact of loss balance factor $\lambda$
Figure 8. The impact of residual weight $\delta$
Figure 9. The impact of dimensionality of time step embedding $d$
In order to achieve a more detailed understanding, we selected some hyperparameters: The loss balance factor $\lambda$, the residual weight $\delta$ and the dimensionality of time step embedding $d$. The results of these experiments are presented in Figure 9-9 by adjusting them within a certain range.
$\bullet$ Loss balance factor $\lambda$. The loss balance factor $r$ affects the performance of RDiffBR, which is achieved by adjusting the emphasis between the recommendation task and the reconstruction task. A high value of $r$ may neglect the core recommendation task and thereby compromise the overall performance.
$\bullet$ Residual weight $\delta$. The appropriate selection of residual weights is crucial for achieving embedding generation, this is because too high or too low values will lead to a decline in performance.
$\bullet$ Dimensionality of time step embedding $d$. As the value of d increases, the model performance initially improves. This is because the time step embedding can flexibly adjust the denoising strategy according to different noise stages. However, when the value of d is excessively large, the input embedding will gradually be weakened during the calculation process, thereby negatively affecting the model performance.
3.5. Case Study
Figure 10. A case study for Youshu under $\rho=-3$. The highlighted items are those before the bundle change.
Figure 11. A case study for Youshu under $\rho=3$. The highlighted items are those before the bundle change.
In this section, we analyze the real data instances as show in Figure 11-11. We randomly select two bundles from Youshu and calculate cosine similarity between item-level bundle embeddings $E^{IL}_{B}$ and all items after B-I change. We listed the top 5 items with the highest similarity scores and their corresponding similarity values. It can be seen that the similarity of items(marked in red) that appeared within the bundle will be higher when only using CrossCBR, resulting in insufficient generalization ability. When using RDiffBR w/o R, the model cannot well capture the main items of the bundle. When using RDiffBR, the model can not only express the main items of the bundle but also introduce other items based on the theme to enhance the generalization ability.
3.6. Training Efficiency Study (RQ3)
In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time(seconds) of using BR model and BR model+RDiffBR on two different scale datasets. We record five consecutive training epochs and average them to obtain the one-epoch training time, as shown in Table 4. It shows that RDiffBR only increased training time about 4%.
Table 4. The statistics of one-epoch training time(seconds) for three backbone models and using RDiffBR on them.
Dataset Youshu NetEase
BGCN 1.46 23.5
BGCN+RDiffBR 1.51 24.5
CrossCBR 1.06 8.7
CrossCBR+RDiffBR 1.11 9.1
BundleGT 22.7 426.2
BundleGT+RDiffBR 23.3 434.8
4. RELATED WORK
4.1. Bundle Recommendation
Bundle recommendation aims to recommend a bundle of items related to a specific theme. Early work employed neural network methods for bundle recommendation (Cao et al., 2017; Chen et al., 2019b; He et al., 2019; Liu et al., 2014). Recently, graph neural networks have become the mainstream technology for bundle recommendation due to its strong ability to encode higher-order information. We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Sun et al., 2024). $\bullet$ Unified learning. The unified representation learning strategy aims to learn representations of users and bundles within a unified framework. There are some existing research studies that utilized this strategy, which integrated users, items and bundles into a unified user-bundle-item graph (Chen et al., 2019a) or hypergraph  (Yu et al., 2022), and propagated neighboring information to get the final representations of users and bundles. $\bullet$ Separate Learning. The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction. This works (Chang et al., 2020; Ke et al., 2023b; Tan et al., 2021; Li et al., 2023; Wei et al., 2023) perform representation learning on several views separately and then fuse the specific predictions of views. $\bullet$ Cooperative Learning. This approach is similar to the separate learning strategy in that it learns user and bundle representations from bundle view and item view but the key difference is that, techniques such as contrastive learning (Zhao et al., 2022; Ma et al., 2022, 2024a; Zou et al., 2024), knowledge distillation (Ke et al., 2023a; Ren et al., 2023) are employed cross views to achieve mutual enhancement.
However, most of bundle recommendation models have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. So they lack the ability to adapt to item-level dynamic variability scenario. It is necessary to study our task for real-world application scenarios.
4.2. Diffusion Models for Recommendation
The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation. Diffusion model plays three main roles in recommendation systems (Lin et al., 2024b): $\bullet$ Diffusion for data augmentation. By generating realistic samples to expand the original training data, such as sequence data (Wu et al., 2023; Liu et al., 2023), multi-field features (Kotelnikov et al., 2023; Zhang et al., 2024), user-item interactions (Liu et al., 2024) and side information (Jiang et al., 2024b, a), diffusion models can enhance the performance of recommendation systems from multiple perspectives. $\bullet$ Diffusion for Representation Enhancement. Diffusion models converts the input data into embeddings to address some key challenges of recommendation. Lots of studies have been proposed such as denoising the implicit user feedback (Zhao et al., 2024), out-of-distribution issue (Zhao et al., 2025) and multi-modal learning (Ma et al., 2024b). $\bullet$ Diffusion as Recommendation Model. When using diffusion models as recommendation models, the input at this stage can be structured datas or embeddings from other encoders. The related research can be classified into three categories: collaborative recommendation (Wang et al., 2023; Hou et al., 2024), context-aware recommendation (Li et al., 2024b; Qin et al., 2024), and other applications (Lin et al., 2024a). $\bullet$ Diffusion for Content Presentation. Diffusion models can further enhance user satisfaction by presenting generated content to target users, rather than conducting item recommendations. Existing research works can be roughly classified into three categories: fashion generation (Xu et al., 2024), advertising creative generation (Yang et al., 2024) and general content generation (Wang et al., 2024).
However, simply applying diffusion model on bundle models is insufficient to address the issue of item-level dynamic variability. There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
5. Conclusion and Future Work
In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, which is a model-agnostic generative framework that can help a BR model adapt this scenario. Given the item-level recommendation model of the bundle, RDiffBR actively injects Gaussian noise into these embeddings and iteratively removes the noise during the reverse process. To guide the reverse process, we designed a residual app module. In the inference stage, we used this module to generate new item-level embeddings. A large number of experimental results show that RDiffBR enables the backbone BR model to achieve significant performance improvement in scenarios with dynamic changes of B-I. Future improvements of RDiffBR may include: 1) allowing RDiffBR to be directly inserted into the well-trained bundle model and take effect; 2) Using diffusion to model bundle features as distridution rather than fixed features that have limitations in B-I dynamic variability scenario.
Acknowledgements.

### v3

@@
 Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation
-Abstract.
+Abstract
 Existing solutions for bundle recommendation (BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles.
@@
 For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments.
-Our empirical study demonstrates that the performance of mainstream BR models will fluctuate or even decline regarding item-level variability.
+Our empirical study demonstrates that the performance of mainstream BR models may fluctuate or decline under item-level variability.
 This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario.
-During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by BR model to represent bundle theme via a forward-reverse process.
+During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by the BR model to represent bundle theme via a forward-reverse process.
 In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings.
-In particular, the residual connection in our residual approximator significantly enhances item-level bundle embeddings generation ability of BR models.
+In particular, the residual connection in our residual approximator significantly enhances BR models’ ability to generate high-quality item-level bundle embeddings.
 Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
-Codes and datasets are available at https://anonymous.4open.science/r/RDiffBR.
-Bundle Recommendation, Diffusion Model, Dynamic Variability Scenario
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; New York,USAisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Recommender systems
-1. INTRODUCTION
+Codehttps://github.com/WUT-IDEA/RDiffBR
+Extended versionhttps://arxiv.org/abs/2507.03280
+Introduction
 With the rapid development of online platforms, bundle recommendation (BR) has become an important tool for enhancing user experience and promoting sales.
@@
 As illustrated in Figure 1, Bundle $1$ initially includes ‘T-shirt’, ‘pants’, ‘shoes’ and ‘watch’ with the theme of ‘casual outfit’.
-Figure 1. An example shows bundle-item affiliation is changed while bundle theme is relatively stable.
-[…]
+Figure 1: An example showing that bundle-item affiliation is changed while bundle theme remains relatively stable.
 In practice, platforms can add the ‘hat’ to align with summer promotions or remove the ‘watch’ based on user feedback.
 An intuitive idea is to regard it as a new bundle.
-However, such item-level changes will not cause bundle theme shifting although bundle theme is jointly determined by the items within the bundle.
+However, such item-level changes may not definitely cause bundle theme shifting although bundle theme is jointly determined by the items within the bundle.
 As shown in Figure 1, the theme of Bundle 2 and Bundle 3 remain ‘casual outfit’ by item variation.
@@
 In recent years, lots of works have focused on bundle recommendation and achieved good results.
-Including factorization models represented by LIRE (Liu et al., 2014), EFM (Cao et al., 2017), attention-based models represented by DAM (Chen et al., 2019b), Attlist (He et al., 2019) and GNN-based models represented by BGCN (Chang et al., 2020), CrossCBR (Ma et al., 2022).
-However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline in B-I dynamic variability scenarios we have mentioned above.
+Including factorization models represented by LIRE (LIRE), EFM (EFM), attention-based models represented by DAM (DAM), Attlist (AttList) and GNN-based models represented by CrossCBR (CrossCBR), BundleGT (BundleGT) and DCBR (DCBR).
+However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline under B-I dynamic variability scenarios we have mentioned above.
 For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases (see Section 3.2 for details).
@@
 Experiments on six bundle models and four public datasets from different domains show that RDiffBR significantly improves the performance of bundle models in dynamic B-I variability scenario.
-The main contributions of this work are summarized as follows:
-$\bullet$ We propose a generative residual diffusion framework for bundle recommendation to address item-level dynamic variability issue.
-$\bullet$ Our residual approximator in diffusion can significantly enhance item-level bundle embeddings generation ability of BR models.
-$\bullet$ Extensive experiments on four datasets from different domains show that RDiffBR enhances the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
-Figure 2. Overview of the proposed RDiffBR framework.
-2. Methodology
-Generally speaking, there is an approache to deal with the item-level dynamic variability.
-Utilize side information (such as item attributes) to introduce additional features but requires architectural modifications to the original model.
-In contrast, our RDiffBR does not require additional information.
-The following will give our technical details.
-2.1. Problem Definition
-$U(|U|=M)$ is denoted as the user set, $B(|B|=N)$as the bundle set, and $I(|I|=O)$ as the item set.
+The main contributions of this work are summarized as follows.
+First, we propose a generative residual diffusion framework for bundle recommendation to address item-level dynamic variability issue.
+Secondly, our residual approximator in diffusion can significantly enhance the item-level bundle embedding generation ability of BR models.
+Lastly, extensive experiments on four datasets from different domains show that RDiffBR enhances the performance of Recall and NDCG of the BR backbone models by up to 23%, while only increasing the training time by about 4%.
+Figure 2: Overview of the proposed RDiffBR framework.
+Related Work
+Bundle Recommendation.
+Bundle recommendation aims to recommend a bundle of items related to a specific theme.
+Early work employed neural network methods for bundle recommendation (EFM; DAM; AttList; LIRE).
+Recently, graph neural networks have become the mainstream technology for bundle recommendation.
+We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Survey_bundle).
+$\bullet$ Unified learning.
+The unified representation learning strategy aims to learn representations of users and bundles within a unified user-bundle-item graph (POG; UHBR).
+$\bullet$ Separate learning.
+The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction (BGCN; IHBR; BundleGT; HBGCN; CMRec).
+$\bullet$ Cooperative learning.
+This approach learns user and bundle representations from bundle view and item view and uses techniques such as contrastive learning (MIDGN; CrossCBR; MultiCBR; HIDGN), knowledge distillation (HyperMBR; DGMAE) or diffusion (DCBR) to achieve cross views mutual enhancement.
+However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability scenario.
+It is necessary to study our task for real-world application scenarios.
+Diffusion Models for Recommendation.
+The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation.
+Diffusion model plays three main roles in recommendation systems (survey_diffusion): $\bullet$ Diffusion for data augmentation.
+This approach generating realistic samples to expand the original training data, such as sequence data (Diff4rec; Plug_diff) and side information (DiffKG; DiffMM).
+$\bullet$ Diffusion for representation enhancement.
+Diffusion models converts the input data into embeddings to address some key challenges of recommendation.
+Lots of studies have been proposed such as denoising the implicit user feedback (DDRM), sequential recommendation (DiffuRec), out-of-distribution issue (CausalDiffRec) and multi-modal learning (MCDRec).
+$\bullet$ Diffusion as recommendation model.
+This approach can be classified into three categories: collaborative recommendation (DRM; CFDiff), context-aware recommendation (RecDiff; Diff-POI), and other applications (DCDR).
+However, a straightforward application of diffusion models on bundle recommendation is insufficient to address the issue of item-level dynamic variability.
+There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
+Methodology
+While it is intuitive to leverage side information—such as item attributes—to enrich the feature space, doing so typically demands non-trivial architectural changes to the original model.
+In contrast, our RDiffBR operates without any external information and leaves the base architecture untouched.
+The technical details are presented below.
+Problem Definition
+$U(|U|=M)$ is denoted as the user set, B(|B|=N)B(|B|=N)as the bundle set, and $I(|I|=O)$ as the item set.
 The user-bundle interactions, user-item interactions, and bundle-item affiliations are represented as $X_{M\times N}=\{(u,b)\mid u\in U,b\in B\}$, $Y_{M\times O}=\{(u,i)\mid u\in U,i\in I\}$ and $Z_{N\times O}=\{(b,i)\mid b\in B,i\in I\}$ with a binary value at each entry respectively.
@@
 Input: The user-bundle interactions $X_{M\times N}$, user-item interactions $Y_{M\times O}$, and bundle-item affiliations $Z_{N\times O}$.
-Output: A well-trained bundle recommendation estimates the probabilty $\hat{y}_{ub}$ of a user $u$ interacting with a bundle $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$.
+Output: A well-trained bundle recommendation estimates the probability $\hat{y}_{ub}\in(0,1)$ of a user $u$ interacting with unseen bundles $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$.
 Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation.
 And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
-2.2. RDiffBR Framework
+RDiffBR Framework
 Considering the successful experience of diffusion as a generative model and the sparsity of input data, our RDiffBR framework introduces a residual diffusion model in the latent space of interactions.
@@
 Since $E^{IL}_{b}$ represents the bundle theme generated by the BR model, we iteratively process $E^{IL}_{b}$ through a forward-backward process, jointly optimizing the BR model and the residual approximator module that we have designed based on the task.
-In the inference stage, RDiffBR applies this optimized forward-reverse process to $E^{IL}_{b}$ from dynamic variability B-I affiliations $Z^{*}$ to obtain new $E^{IL}_{b}$ for bundle recommendation, enhancing BR model’s ability for this scenario.
-2.2.1. Residual Diffusion.
+In the inference stage, RDiffBR applies this optimized forward-reverse process to $E^{IL}_{b}$ from dynamic variability B-I affiliations $Z^{*}$ to obtain new $E^{IL}_{b}$, enhancing BR model’s ability for this scenario.
+Residual Diffusion
 $\bullet$ Forward process.
@@
 As shown in the bottom of residual diffusion module in Figure 2, Gaussian noises are continuously incorporated into $e_{0}$ with adjustable scales and steps:
-(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}).$
-Based on the reparameterization (Ho et al., 2020) and the additivity of independent Gaussian noises (Luo, 2022), we can directly obtain $e_{t}$ from $e_{0}$.
-(2) $q(e_{t}|e_{0})=\mathcal{N}(e_{t},\sqrt{\bar{\alpha}_{t}}e_{0},(1-\bar{\alpha_{t}})\boldsymbol{I}),$
+$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}).$ (1)
+Based on the reparameterization (Diffusion) and the additivity of independent Gaussian noises, we can directly obtain $e_{t}$ from $e_{0}$.
+$q(e_{t}|e_{0})=\mathcal{N}(e_{t},\sqrt{\bar{\alpha}_{t}}e_{0},(1-\bar{\alpha_{t}})\boldsymbol{I}),$ (2)
 where $\bar{\alpha_{t}}=\prod_{t=1}^{t}\alpha_{t},\alpha_{t}=1-\beta_{t},\beta_{t}\in(0,1)$ controls the Gaussian noise scales added at each step $t$.
-To regulate the added noises in $e_{1:T}$, we follow  (Wang et al., 2023) using a linear variance noise schedule in the forward process:
-(3) $1-\bar{\alpha}_{t}=s\cdot\left[\alpha_{\min}+\frac{t-1}{T-1}(\alpha_{\max}-\alpha_{\min})\right],$
-where T is the total forward step and s $\in$ (0, 1) controls the noise scale. The condition $\alpha_{min}$ $<$ $\alpha_{max}$in (0, 1) indicates the upper and lower limits of the added noise.
+To regulate the added noises in $e_{1:T}$, we follow  (DRM) using a linear variance noise schedule in the forward process:
+$1-\bar{\alpha}_{t}=s\cdot\left[\alpha_{\min}+\frac{t-1}{T-1}(\alpha_{\max}-\alpha_{\min})\right],$ (3)
+where T is the total forward step and s $\in$ (0, 1) controls the noise scale. The condition $\alpha_{min}$ << $\alpha_{max}$ $\in$ (0, 1) indicates the upper and lower limits of the added noise.
 $\bullet$ Reverse process.
@@
 In each reverse step, a residual approximator module to denoise them.
-(4) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\Sigma_{\theta}(\hat{e}_{t},t)),$
+$p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\Sigma_{\theta}(\hat{e}_{t},t)),$ (4)
 where $\hat{e}_{t}$ is denoised embedding in the reverse step t, $\theta$ and are the learnable parameters of residual approximator.
-This module are executed iteratively in the reverse process until the reconstruction of final embedding $\hat{e}_{0}$
+This module is executed iteratively in the reverse process until reconstructing final embedding $\hat{e}_{0}$
 $\bullet$ Residual approximator.
@@
 Through Bayes’ theorem, we can derive:
-(5) $q(e_{t-1}|e_{t},e_{0})\propto\mathcal{N}(e_{t-1};\tilde{\mu}(e_{t},e_{0},t),\sigma^{2}(t)\boldsymbol{I}),\mathrm{~{}where}$
-$\begin{cases}\tilde{\mu}(e_{t},e_{0},t)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}e_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}e_{0}\\ \sigma^{2}(t)=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},&\end{cases}$
-$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (Zhao et al., 2024) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
-(6) $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~{}where}$
+$q(e_{t-1}|e_{t},e_{0})\propto\mathcal{N}(e_{t-1};\tilde{\mu}(e_{t},e_{0},t),\sigma^{2}(t)\boldsymbol{I}),\mathrm{~where}$ (5)
+[…]
+$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
+$p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~where}$ (6)
 $\mu_{\theta}(\hat{e}_{t},t)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\hat{e}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\tilde{e}_{0},$
@@
 Specifically, for embedding $e_{t}$ in the reverse step t:
-(7) $\tilde{e}_{0}=\delta f_{\theta}(\hat{e}_{t},t)+(1-\delta)e_{t^{\prime}},$
+$\tilde{e}_{0}=\delta f_{\theta}(\hat{e}_{t},t)+(1-\delta)e_{t^{\prime}},$ (7)
 where $\tilde{e}_{0}$ is the predicted value of $e_{0}$ by residual approximator with parameter $\theta$.
@@
 $e_{t^{\prime}}$ is noisy embedding in the forward step $t^{\prime}$.
-Step information t is encoded by sinusoidal positional encoding (Ho et al., 2020), and these two inputs are concated as input.
-2.2.2. Loss Function.
+Step information t is encoded by sinusoidal positional encoding (Diffusion), and these two inputs are concated as input.
+Loss Function.
 To optimize the forward-reverse process, it is necessary to minimize the variational lower bound of the predicted embedding.
 According to the KL divergence based on the multivariate Gaussian distribution, the reconstruction loss of within one training iteration can be expressed as:
-(8) $\mathcal{L}_{r}$ $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{e}_{t-1}|\mathbf{e}_{t},\mathbf{e}_{0}\right)\|p_{\theta}\left(\mathbf{e}_{t-1}|\mathbf{e}_{t}\right)\right)\right]$
+$\mathcal{L}_{r}$ $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{e}_{t-1}|\mathbf{e}_{t},\mathbf{e}_{0}\right)\|p_{\theta}\left(\mathbf{e}_{t-1}|\mathbf{e}_{t}\right)\right)\right]$ (8)
 $=\mathbb{E}_{q(\mathbf{e}_{t}|\mathbf{e}_{0})}\left[\frac{1}{2\sigma^{2}(t)}\left[\|\mu_{\theta}(\mathbf{e}_{t},t)-\mu(\mathbf{e}_{t},\mathbf{e}_{0},t)\|_{2}^{2}\right]\right].$
-Furthermore, by substituting Eq.(5) and Eq.(6) into Eq.(8) and simplifying (Wang et al., 2023) we can drive:
-(9) $\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{t\sim\mathcal{U}(1,T)}\mathbb{E}_{q}\left[||e_{0}-\hat{e}_{\theta}(e_{t},e_{t^{\prime}},t)||_{2}^{2}\right],$
+Furthermore, by substituting Eq.(5) and Eq.(6) into Eq.(8) and simplifying (DRM) we can drive:
+[…]
 where $\mathcal{U}$ means uniformly sampling t from $\{1,2,...,T\}$.
@@
 We have designed a loss balancing factor $\lambda$ to adjust the weight of reconstruction loss and jointly train the bundle model and residual approximator together:
-(10) $\mathcal{L}=\mathcal{L}_{BR\mathrm{~{}model}}+\lambda\mathcal{L}_{\mathrm{diff}}.$
+$\mathcal{L}=\mathcal{L}_{BR\mathrm{~model}}+\lambda\mathcal{L}_{\mathrm{diff}}.$ (10)
 The training step of residual approximator is illustrated in Algorithm 1.
 Specifically, Lines 2-5 represent the forward process, Line 6 describes the reverse process using residual approximator, and Lines 7-8 address optimization.
+1
 Input: interaction data $X$, item-level bundle embedding $e_{0}$, diffusion step $T$, randomly initialized residual approximator MLP $f_{\theta}$.
@@
 3   Sample a batch of U-B interactions $\bar{X}\subset X$.
-3
+4
 5 for $all\ (u,b)\in\bar{X}$ do
 6      Sample $t\sim\mathcal{U}(1,T)$;
-6
+7
 8     Compute $e_{t}$ given $e_{0}$ and $t$ via $q(e_{t}|e_{0})$ in Eq.(2);
-8
+9
 10     Reconstruct $\tilde{e}_{0}$ through $f_{\theta}$ in Eq.(7);
 11     Calculate $\mathcal{L}$ by Eq.(10);
-11
+12
 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
 14   end for
-14
-15until converged;
+15
+16until converged;
 Output: optimized $\theta.$
 Algorithm 1 Residual approximator training process
-2.2.3. Inference.
+Inference.
 In the inference stage, we take the item-level bundle embedding $e^{*}_{0}$ in the B-I dynamic variability scenario as input.
 Subsequently, RDiffBR continuously introduces noise into $e^{*}_{0}$ in the sequence $e^{*}_{0}\to e^{*}_{1}\to...\to e^{*}_{T^{\prime}}$ for $T^{\prime}$ steps in the forward process and then we set $\hat{e}_{{T^{\prime}}}=e^{*}_{T^{\prime}}$ to execute reverse process $\hat{e_{T^{\prime}}}\to\hat{e}_{{T^{\prime}}-1}\to...\to\hat{e}_{0}$ to generate new item-level bundle embedding $\hat{e}_{0}$.
-The reverse ignores the variance and uses $\hat{e}_{t-1}=\mu_{\theta}(\hat{e}_{t},e^{*}_{0},t)$ for directly inference following (Liang et al., 2018).
+The reverse ignores the variance and uses $\hat{e}_{t-1}=\mu_{\theta}(\hat{e}_{t},e^{*}_{0},t)$ for directly inference following (MultiVAE).
 We replace the original $e^{*}_{0}$ with $\hat{e}_{0}$, and then conduct bundle recommendation following the original way of the bundle model.
-Finally enhancing the ranking of bundles in dynamic variation scenarios.
+Finally the ranking of bundles is enhanced.
 The inference procedure of residual approximator is illustrated in Algorithm 2.
@@
 Algorithm 2 Inference process
-3. EXPERIMENTS
-Our experiments aim to answer the following research questions:
-$\bullet$ RQ1: How does RDiffBR perform on different datasets of backbone bundle models under various degrees of B-I variability?
-$\bullet$ RQ2: What is the contributions of residual connection in RDiffBR to its overall performance?
+Experiments
+Our experiments aim to answer the following questions:
+$\bullet$ RQ1: How does RDiffBR perform with backbone bundle models under various degrees of B-I variability?
+$\bullet$ RQ2: What is the contribution of residual connection in RDiffBR to its overall performance?
 $\bullet$ RQ3: How does the efficiency of RDiffBR?
-3.1. Experimental Settings
+Experimental Settings
 $\bullet$ Datasets.
 Four publicly bundle datasets from different domains are summarized in Table 1.
-Youshu (Chen et al., 2019b) for book list recommendation; NetEase (Cao et al., 2017) for music playlist recommendation; iFashion (Chen et al., 2019a) for fashion outfit recommendation; In addition, we introduce another meal recommendation dataset $\mathrm{MealRec+}$ (Li et al., 2024a), where the meal consisted of individual course items is treated as bundle.
-Table 1. Dataset Statistics.
-Dataset #U #I #B #U-I #U-B #B-I #Avg.I/B #Dens.B-I
-Youshu 8,039 32,770 4,771 138,515 51,377 176,667 37.03 0.11%
-NetEase 18,528 123,628 22,864 1,128,065 302,303 1,778,838 77.8 0.06%
-iFashion 53,897 42,563 27,694 2,290,645 1,679,708 106,916 3.86 0.01%
-$\mathrm{MealRec+}$ 1,575 7,280 3,817 151,148 46,767 11,451 3 0.77%
-$\bullet$ Baselines.
-We implement RDiffBR on six representative BR models with official code released by their authors.
-BGCN (Chang et al., 2020) utilizes graph convolutional networks to separately learn the representations of bundle view and item view.
-HyperMBR (Ke et al., 2023a) models two views in hyperbolic space to learn their accurate representations.
-MIDGN (Zhao et al., 2022) refines the user-item/bundle-item graph of different intents from multiple views.
-CrossCBR (Ma et al., 2022) improves two separate views by constrastive learning.
-MultiCBR (Ma et al., 2024a) first fuses the multi-view representations into a unified one, and then adopts a contrastive loss on the unified representations.
-BundleGT (Wei et al., 2023) learns the strategy-aware user and bundle representation in bundle recommendation.
-There are still some earlier research outcomes (for instance, BundleNet (Deng et al., 2020)) that have been studied to be less effective than the aforementioned methods.
-Therefore, we do not take them into consideration.
-$\bullet$ Experimental process.
-According to previous works (Li et al., 2024a) (Ma et al., 2022), the training, validation and testing sets of user-bundle interactions $X$ are randomly splited with specific ratio.
+Youshu (DAM) for book list recommendation; NetEase (EFM) for music playlist recommendation; iFashion (POG) for fashion outfit recommendation; In addition, we introduce another meal recommendation dataset $\mathrm{MealRec+}$ (mealrec+), where the meal consisted of individual course items is treated as bundle.
+[…]
+# U 8,039 18,528 53,897 1,575
+# I 32,770 123,628 42,563 7,280
+# B 4,771 22,864 27,694 3,817
+# U-I 138,515 1,128,065 2,290,645 151,148
+# U-B 51,377 302,303 1,679,708 46,767
+# B-I 176,667 1,778,838 106,916 11,451
+# Avg. I/B 37.03 77.8 3.86 3
+Dens.B-I 0.11% 0.06% 0.01% 0.77%
+Table 1: Statistics of datasets.
+Dataset Youshu NetEase
+$\rho$ -4 -2 0 2 4 -4 -2 0 2 4
+Metrics R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20
+BGCN .1897 .1047 .2254 .1278 .2367 .1380 .2214 .1247 .2138 .1204 .0581 .0309 .0634 .0333 .0643 .0339 .0646 .0341 .0622 .0332
++RDiffBR .2178 .1278 .2391 .1415 .2453 .1466 .2470 .1423 .2407 .1396 .0614 .0321 .0657 .0340 .0662 .0343 .0668 .0345 .0649 .0340
+%improv. 14.83 22.15 6.06 10.74 3.61 6.25 11.57 14.11 12.60 16.03 5.56 3.98 3.63 1.99 2.89 1.14 3.47 1.07 4.42 2.54
+MIDGN .2388 .1373 .2557 .1472 .2598 .1496 .2578 .1495 .2566 .1489 .0540 .0278 .0625 .0330 .0636 .0337 .0639 .0337 .0640 .0337
++RDiffBR .2498 .1441 .2621 .1498 .2645 .1515 .2652 .1517 .2643 .1516 .0568 .0293 .0649 .0340 .0661 .0347 .0667 .0348 .0672 .0350
+%improv. 4.62 4.98 2.53 1.72 1.79 1.32 2.87 1.49 3.00 1.82 5.21 5.27 3.89 2.91 3.96 2.92 4.45 3.25 4.95 4.14
+HyperMBR .1464 .0953 .2126 .1290 .2649 .1581 .2167 .1285 .1897 .1162 .0588 .0313 .0675 .0354 .0717 .0377 .0689 .0363 .0661 .0345
++RDiffBR .1789 .1109 .2229 .1327 .2669 .1594 .2280 .1347 .2128 .1275 .0627 .0332 .0694 .0369 .0739 .0386 .0704 .0369 .0667 .0350
+%improv. 22.20 16.42 4.85 2.87 0.75 0.83 5.20 4.87 12.14 9.75 6.67 5.97 2.88 4.41 3.07 2.51 2.25 1.79 0.85 1.50
+CrossCBR .2677 .1584 .2763 .1626 .2776 .1640 .2788 .1649 .2800 .1656 .0715 .0386 .0776 .0419 .0794 .0431 .0802 .0436 .0798 .0434
++RDiffBR .2744 .1640 .2814 .1659 .2827 .1663 .2816 .1664 .2816 .1662 .0772 .0419 .0818 .0436 .0826 .0443 .0829 .0444 .0826 .0445
+%improv. 2.52 3.57 1.84 2.02 1.85 1.42 1.01 0.92 0.56 0.40 7.99 8.50 5.36 4.04 4.09 2.67 3.43 1.70 3.50 2.51
+MultiCBR .2770 .1649 .2769 .1649 .2791 .1652 .2767 .1641 .2770 .1648 .0843 .0463 .0867 .0472 .0876 .0475 .0880 .0479 .0879 .0479
++RDiffBR .2831 .1674 .2840 .1677 .2824 .1673 .2825 .1669 .2842 .1675 .0859 .0469 .0879 .0478 .0881 .0480 .0883 .0482 .0893 .0487
+%improv. 2.20 1.55 2.56 1.71 1.19 1.25 2.10 1.71 2.60 1.61 1.91 1.32 1.37 1.27 0.57 1.18 0.40 0.75 1.60 1.52
+BundleGT .2687 .1590 .2872 .1683 .2877 .1694 .2873 .1697 .2837 .1679 .0729 .0367 .0869 .0465 .0874 .0468 .0874 .0468 .0871 .0465
++RDiffBR .2829 .1653 .2927 .1708 .2922 .1707 .2939 .1712 .2903 .1702 .0759 .0384 .0900 .0479 .0896 .0479 .0897 .0477 .0895 .0476
+%improv. 5.27 3.96 1.90 1.49 1.57 0.79 2.30 0.93 2.31 1.40 4.06 4.75 3.49 3.16 2.52 2.42 2.67 1.86 2.76 2.37
+[…]
+$\rho$ -4 -2 0 2 4 -4 -2 0 2 4
+Methods R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20
+BGCN .0473 .0331 .0560 .0395 .0602 .0426 .0607 .0428 .0604 .0427 .1068 .1888 .1213 .2094 .1325 .2247 .1326 .2250 .1375 .2323
++RDiffBR .0488 .0341 .0568 .0398 .0606 .0428 .0608 .0429 .0607 .0429 .1180 .2078 .1293 .2180 .1408 .2401 .1387 .2371 .1388 .2384
+%improv. 3.09 2.91 1.55 0.58 0.61 0.45 0.16 0.23 0.57 0.47 10.07 10.55 4.08 6.53 6.82 6.26 5.38 4.57 2.63 0.92
+MIDGN .0372 .0257 .0549 .0380 .0672 .0482 .0428 .0279 .0441 .0294 .0820 .1564 .1017 .1816 .1121 .1908 .1129 .1919 .1115 .1880
++RDiffBR .0426 .0290 .0628 .0437 .0763 .0553 .0534 .0348 .0545 .0363 .0826 .1604 .1075 .1930 .1179 .2050 .1145 .2012 .1146 .2027
+%improv. 14.55 12.83 14.36 14.87 13.53 14.65 24.66 24.83 23.48 23.77 2.60 0.82 6.29 5.69 7.44 5.13 4.87 1.48 7.82 2.79
+HyperMBR .0448 .0336 .0687 .0494 .0815 .0580 .0773 .0545 .0744 .0521 .0859 .1498 .0939 .1595 .1069 .1778 .1057 .1781 .1106 .1887
++RDiffBR .0462 .0349 .0692 .0504 .0830 .0595 .0797 .0567 .0773 .0550 .0874 .1585 .0995 .1760 .1122 .1944 .1205 .1988 .1247 .2144
+%improv. 3.21 4.09 0.81 2.01 1.75 2.69 3.01 4.01 3.92 5.42 5.82 1.70 10.33 5.90 9.34 4.94 11.61 14.00 13.62 12.82
+CrossCBR .0624 .0491 .0877 .0683 .0999 .0772 .1046 .0799 .1081 .0831 .0612 .1048 .0819 .1417 .0938 .1688 .1084 .2041 .1504 .2599
++RDiffBR .0720 .0554 .0928 .0716 .1027 .0794 .1060 .0812 .1088 .0834 .0681 .1203 .0926 .1531 .1026 .1788 .1241 .2243 .1611 .2669
+%improv. 15.36 12.87 5.75 4.78 2.83 2.82 1.31 1.64 0.60 0.32 14.80 11.40 8.01 13.15 5.95 9.37 9.89 14.48 2.69 7.12
+MultiCBR .0672 .0588 .1080 .0905 .1297 .1064 .1343 .1101 .1381 .1133 .0989 .1782 .1147 .1841 .1305 .2092 .1476 .2376 .1625 .2592
++RDiffBR .0694 .0608 .1094 .0916 .1313 .1080 .1350 .1111 .1385 .1139 .1066 .1889 .1217 .1945 .1375 .2164 .1524 .2436 .1633 .2598
+%improv. 3.29 3.31 1.23 1.16 1.20 1.47 0.53 0.97 0.27 0.58 6.04 7.83 5.63 6.10 3.44 5.40 2.51 3.19 0.25 0.50
+BundleGT .0444 .0352 .0690 .0517 .0954 .0724 .0899 .0653 .0960 .0707 .0732 .1281 .0923 .1556 .1093 .1761 .1217 .2134 .1579 .2585
++RDiffBR .0456 .0361 .0701 .0530 .0970 .0740 .0909 .0660 .0970 .0712 .0863 .1512 .1082 .1770 .1315 .2077 .1414 .2346 .1707 .2733
+%improv. 2.68 2.58 1.64 2.34 1.61 2.10 1.02 1.04 1.10 0.71 18.03 17.88 13.75 17.20 17.97 20.35 9.91 16.21 5.73 8.05
+[…]
+$\bullet$ BR Backbones.
+We implement RDiffBR on six representative BR models with official codes released by their authors.
+BGCN (BGCN) utilizes graph convolutional networks to separately learn the representations of bundle view and item view.
+HyperMBR (HyperMBR) models two views in hyperbolic space to learn their accurate representations.
+MIDGN (MIDGN) refines the user-item/bundle-item graph of different intents from multiple views.
+CrossCBR (CrossCBR) improves two separate views by constrastive learning.
+MultiCBR (MultiCBR) first fuses the multi-view representations into a unified one, and then adopts a contrastive loss on the unified representations.
+BundleGT (BundleGT) learns the strategy-aware user and bundle representation in bundle recommendation.
+Figure 3: The processing of the B-I affiliation to match the scenario of B-I dynamic variability.
+$\bullet$ Experimental Preprocess and Metrics.
+According to previous works (CrossCBR) (mealrec+), the training, validation and testing sets of user-bundle interactions $X$ are randomly splited with specific ratio.
 It is not feasible to simply add new bundles in bundle-item affiliations $Z$ under the B-I variability assumptions.
@@
 In testing phase, we design $Z_{\mathrm{test}}$ by adjusting $Z_{\mathrm{train\&val}}$ with adding or removing subsets:
-(11) $Z_{\mathrm{test}}=\begin{cases}Z_{\mathrm{train\&val}}\cup\left(\mathcal{\cup}_{i=6}^{5+\rho}\mathcal{E}_{i}\right)&\mathrm{if}~{}\rho>0\\ Z_{\mathrm{{train\&val}}}&\mathrm{if~{}\rho=0}\\ Z_{\mathrm{train\&val}}\setminus\left(\mathcal{\cup}_{i=6-|\rho|)}^{5}\mathcal{E}_{i}\right)&\mathrm{if~{}\rho<0}\end{cases}$
+$Z_{\mathrm{test}}=\begin{cases}Z_{\mathrm{train\&val}}\cup\left(\mathcal{\cup}_{i=6}^{5+\rho}\mathcal{E}_{i}\right)&\mathrm{if}~\rho>0\\ Z_{\mathrm{{train\&val}}}&\mathrm{if~\rho=0}\\ Z_{\mathrm{train\&val}}\setminus\left(\mathcal{\cup}_{i=6-|\rho|)}^{5}\mathcal{E}_{i}\right)&\mathrm{if~\rho<0}\end{cases}$ (11)
 Where $\cup$ indicates set union and $\setminus$ indicates set difference.
@@
 $\rho>0$ indicates adding set $\Delta^{+}$ to $Z_{\mathrm{test}}$, $\rho<0$ indicates removing set $\Delta^{-}$ from $Z_{\mathrm{test}}$ and $\rho=0$ indicates $Z_{\mathrm{test}}$ invariability.
-The well-trained bundle models predict user-bundle interactions $X$ on $Z_{test}$.
+The well-trained bundle models predict user-bundle interactions $X$ on $Z_{\mathrm{test}}$.
 The processing of the B-I affiliation is shown in Figure 3.
-In this way, we can control the degree of B-I affiliaitons variation and simulate the scenario of B-I dynamic variability.
-Figure 3. The processing of the B-I affiliation to match the scenario of B-I dynamic variability.
-$\bullet$ Hyperparameter Settings We set the hyperparameters of six backbone models from their authors and optimize RDiffBR with the Adam optimizer with a learning rate of 0.001.
+In this way, we can control the degree of B-I affiliatons variation and simulate the scenario of B-I dynamic variability.
+Recall@K and NDCG@K are used as the evaluation metrics, where K $\in\{10,20,40,80\}$.
+$\bullet$ Hyperparameter Settings The hyperparameters of six backbone models are set by following their publications and RDiffBR is optimized with the Adam optimizer with a learning rate of 0.001.
 The time steps $T$ and sampling steps $T^{\prime}$ are set to values ranging from [1,200].
@@
 The number of MLP layers is selected within 1, 2, 4, 8.
-The hidden size of the MLP layers are chosen from [8,4096].
+Its hidden sizes are from [8,4096].
 Details of other hyperparameters can be found in our released code.
-All the models are trained using two GTX 3090.
-3.2. Overall Performance (RQ1)
-Table 2.
-Performance comparison between backbone models and our RDiffBR on four datasets($\rho$=-4, -2, 0, 2, 4), while $\rho<0$ indicates $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\setminus\Delta^{-}$, $\rho>0$ indicates adding set $Z_{\mathbf{test}}=Z_{\mathbf{train\&val}}\cup\Delta^{+}$, and $\rho=0$ indicates $Z_{\mathbf{train\&val}}$ invariability.
-Improv.
-indicates the relative improvement of a backbone model+RDiffBR over the itself.
-Dataset Youshu NetEase
-$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
-Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
-BGCN 0.1897 0.1047 0.2254 0.1278 0.2367 0.1380 0.2214 0.1247 0.2138 0.1204 0.0581 0.0309 0.0634 0.0333 0.0643 0.0339 0.0646 0.0341 0.0622 0.0332
-+RDiffBR 0.2178 0.1278 0.2391 0.1415 0.2453 0.1466 0.2470 0.1423 0.2407 0.1396 0.0614 0.0321 0.0657 0.0340 0.0662 0.0343 0.0668 0.0345 0.0649 0.0340
-improv. 14.83% 22.15% 6.06% 10.74% 3.61% 6.25% 11.57% 14.11% 12.60% 16.03% 5.56% 3.98% 3.63% 1.99% 2.89% 1.14% 3.47% 1.07% 4.42% 2.54%
-MIDGN 0.2388 0.1373 0.2557 0.1472 0.2598 0.1496 0.2578 0.1495 0.2566 0.1489 0.0540 0.0278 0.0625 0.0330 0.0636 0.0337 0.0639 0.0337 0.0640 0.0337
-+RDiffBR 0.2498 0.1441 0.2621 0.1498 0.2645 0.1515 0.2652 0.1517 0.2643 0.1516 0.0568 0.0293 0.0649 0.0340 0.0661 0.0347 0.0667 0.0348 0.0672 0.0350
-improv. 4.62% 4.98% 2.53% 1.72% 1.79% 1.32% 2.87% 1.49% 3.00% 1.82% 5.21% 5.27% 3.89% 2.91% 3.96% 2.92% 4.45% 3.25% 4.95% 4.14%
-HyperMBR 0.1464 0.0953 0.2126 0.1290 0.2649 0.1581 0.2167 0.1285 0.1897 0.1162 0.0588 0.0313 0.0675 0.0354 0.0717 0.0377 0.0689 0.0363 0.0661 0.0345
-+RDiffBR 0.1789 0.1109 0.2229 0.1327 0.2669 0.1594 0.2280 0.1347 0.2128 0.1275 0.0627 0.0332 0.0694 0.0369 0.0739 0.0386 0.0704 0.0369 0.0667 0.0350
-improv. 22.20% 16.42% 4.85% 2.87% 0.75% 0.83% 5.20% 4.87% 12.14% 9.75% 6.67% 5.97% 2.88% 4.41% 3.07% 2.51% 2.25% 1.79% 0.85% 1.50%
-CrossCBR 0.2677 0.1584 0.2763 0.1626 0.2776 0.1640 0.2788 0.1649 0.2800 0.1656 0.0715 0.0386 0.0776 0.0419 0.0794 0.0431 0.0802 0.0436 0.0798 0.0434
-+RDiffBR 0.2744 0.1640 0.2814 0.1659 0.2827 0.1663 0.2816 0.1664 0.2816 0.1662 0.0772 0.0419 0.0818 0.0436 0.0826 0.0443 0.0829 0.0444 0.0826 0.0445
-improv. 2.52% 3.57% 1.84% 2.02% 1.85% 1.42% 1.01% 0.92% 0.56% 0.40% 7.99% 8.50% 5.36% 4.04% 4.09% 2.67% 3.43% 1.70% 3.50% 2.51%
-MultiCBR 0.2770 0.1649 0.2769 0.1649 0.2791 0.1652 0.2767 0.1641 0.2770 0.1648 0.0843 0.0463 0.0867 0.0472 0.0876 0.0475 0.0880 0.0479 0.0879 0.0479
-+RDiffBR 0.2831 0.1674 0.2840 0.1677 0.2824 0.1673 0.2825 0.1669 0.2842 0.1675 0.0859 0.0469 0.0879 0.0478 0.0881 0.0480 0.0883 0.0482 0.0893 0.0487
-improv. 2.20% 1.55% 2.56% 1.71% 1.19% 1.25% 2.10% 1.71% 2.60% 1.61% 1.91% 1.32% 1.37% 1.27% 0.57% 1.18% 0.40% 0.75% 1.60% 1.52%
-BundleGT 0.2687 0.1590 0.2872 0.1683 0.2877 0.1694 0.2873 0.1697 0.2837 0.1679 0.0729 0.0367 0.0869 0.0465 0.0874 0.0468 0.0874 0.0468 0.0871 0.0465
-+RDiffBR 0.2829 0.1653 0.2927 0.1708 0.2922 0.1707 0.2939 0.1712 0.2903 0.1702 0.0759 0.0384 0.0900 0.0479 0.0896 0.0479 0.0897 0.0477 0.0895 0.0476
-improv. 5.27% 3.96% 1.90% 1.49% 1.57% 0.79% 2.30% 0.93% 2.31% 1.40% 4.06% 4.75% 3.49% 3.16% 2.52% 2.42% 2.67% 1.86% 2.76% 2.37%
-Dataset iFashion $\mathbf{{MealRec+}}$
-$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
-Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
-BGCN 0.0473 0.0331 0.0560 0.0395 0.0602 0.0426 0.0607 0.0428 0.0604 0.0427 0.1068 0.1888 0.1213 0.2094 0.1325 0.2247 0.1326 0.2250 0.1375 0.2323
-+RDiffBR 0.0488 0.0341 0.0568 0.0398 0.0606 0.0428 0.0608 0.0429 0.0607 0.0429 0.1180 0.2078 0.1293 0.2180 0.1408 0.2401 0.1387 0.2371 0.1388 0.2384
-improv. 3.09% 2.91% 1.55% 0.58% 0.61% 0.45% 0.16% 0.23% 0.57% 0.47% 10.07% 10.55% 4.08% 6.53% 6.82% 6.26% 5.38% 4.57% 2.63% 0.92%
-MIDGN 0.0372 0.0257 0.0549 0.0380 0.0672 0.0482 0.0428 0.0279 0.0441 0.0294 0.0820 0.1564 0.1017 0.1816 0.1121 0.1908 0.1129 0.1919 0.1115 0.1880
-+RDiffBR 0.0426 0.0290 0.0628 0.0437 0.0763 0.0553 0.0534 0.0348 0.0545 0.0363 0.0826 0.1604 0.1075 0.1930 0.1179 0.2050 0.1145 0.2012 0.1146 0.2027
-improv. 14.55% 12.83% 14.36% 14.87% 13.53% 14.65% 24.66% 24.83% 23.48% 23.77% 2.60% 0.82% 6.29% 5.69% 7.44% 5.13% 4.87% 1.48% 7.82% 2.79%
-HyperMBR 0.0448 0.0336 0.0687 0.0494 0.0815 0.0580 0.0773 0.0545 0.0744 0.0521 0.0859 0.1498 0.0939 0.1595 0.1069 0.1778 0.1057 0.1781 0.1106 0.1887
-+RDiffBR 0.0462 0.0349 0.0692 0.0504 0.0830 0.0595 0.0797 0.0567 0.0773 0.0550 0.0874 0.1585 0.0995 0.1760 0.1122 0.1944 0.1205 0.1988 0.1247 0.2144
-improv. 3.21% 4.09% 0.81% 2.01% 1.75% 2.69% 3.01% 4.01% 3.92% 5.42% 5.82% 1.70% 10.33% 5.90% 9.34% 4.94% 11.61% 14.00% 13.62% 12.82%
-CrossCBR 0.0624 0.0491 0.0877 0.0683 0.0999 0.0772 0.1046 0.0799 0.1081 0.0831 0.0612 0.1048 0.0819 0.1417 0.0938 0.1688 0.1084 0.2041 0.1504 0.2599
-+RDiffBR 0.0720 0.0554 0.0928 0.0716 0.1027 0.0794 0.1060 0.0812 0.1088 0.0834 0.0681 0.1203 0.0926 0.1531 0.1026 0.1788 0.1241 0.2243 0.1611 0.2669
-improv. 15.36% 12.87% 5.75% 4.78% 2.83% 2.82% 1.31% 1.64% 0.60% 0.32% 14.80% 11.40% 8.01% 13.15% 5.95% 9.37% 9.89% 14.48% 2.69% 7.12%
-MultiCBR 0.0672 0.0588 0.1080 0.0905 0.1297 0.1064 0.1343 0.1101 0.1381 0.1133 0.0989 0.1782 0.1147 0.1841 0.1305 0.2092 0.1476 0.2376 0.1625 0.2592
-+RDiffBR 0.0694 0.0608 0.1094 0.0916 0.1313 0.1080 0.1350 0.1111 0.1385 0.1139 0.1066 0.1889 0.1217 0.1945 0.1375 0.2164 0.1524 0.2436 0.1633 0.2598
-improv. 3.29% 3.31% 1.23% 1.16% 1.20% 1.47% 0.53% 0.97% 0.27% 0.58% 6.04% 7.83% 5.63% 6.10% 3.44% 5.40% 2.51% 3.19% 0.25% 0.50%
-BundleGT 0.0444 0.0352 0.0690 0.0517 0.0954 0.0724 0.0899 0.0653 0.0960 0.0707 0.0732 0.1281 0.0923 0.1556 0.1093 0.1761 0.1217 0.2134 0.1579 0.2585
-+RDiffBR 0.0456 0.0361 0.0701 0.0530 0.0970 0.0740 0.0909 0.0660 0.0970 0.0712 0.0863 0.1512 0.1082 0.1770 0.1315 0.2077 0.1414 0.2346 0.1707 0.2733
-improv. 2.68% 2.58% 1.64% 2.34% 1.61% 2.10% 1.02% 1.04% 1.10% 0.71% 18.03% 17.88% 13.75% 17.20% 17.97% 20.35% 9.91% 16.21% 5.73% 8.05%
-We present the results of Recall@20 and NDCG@20 for five levels of B-I variations in Table 2.
-We show the results of Recall@40 and NDCG@40 on two datasets for ten levels of B-I variations in Figure 4.
-The following observations can be made:
-$\bullet$ The perfomance of BR models when meeting item-level dynamic variability.
+All the models are trained using two GTX 4090.
+[…]
+Overall Performance (RQ1)
+The following observations are from Table 2.
+$\bullet$ The performance of BR backbones when meeting item-level dynamic variability.
 When $\rho<0$ ($\mathbf{Z_{test}}$ removes $\Delta^{-}$), the performance of bundle models is generally lower than that under $\rho=0$ in most cases.
@@
 This remarkable performance improvement can be attributed to RDiffBR’s generation capability, which reconstructs item-level bundle embeddings in B-I change scenarios to better align with bundle themes rather than overfitting to the B-I affiliations observed during training.
-Figure 4. Performance comparisons of the backbone models and our RDiffBR under different levels of B-I variations in terms of Recall@40 and NDCG@40.
-3.3. Ablation Study (RQ2)
+Ablation Study (RQ2)
 To assess whether residual connections contribute to improving model performance, we set a variant RDiffBR w/o R that omits the residual connection in Eq.(7) and conduct a comprehensive ablation study on three BR models across three datasets, with the results shown in Table 3.
-Table 3. Ablation study where bold indicates the best method, and underlined indicates the second best method($\rho=-3,3$).
-Method Youshu NetEase $\mathrm{MealRec+}$
+[…]
 R@20 N@20 R@20 N@20 R@20 N@20
@@
 BundleGT+RDiffBR w/o R 0.2519 0.1454 0.0520 0.0267 0.2313 0.1416
-Method Youshu NetEase $\mathrm{MealRec+}$
+[…]
 R@20 N@20 R@20 N@20 R@20 N@20
@@
 BundleGT+RDiffBR w/o R 0.2519 0.1454 0.0520 0.0267 0.2312 $\underline{0.1417}$
-Figure 5. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=-3$.
-Figure 6. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=3$.
+Table 3: Ablation study where bold indicates the best method, and underlined indicates the second best method($\rho=-3,3$).
+Figure 5: Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=-3$.
+Figure 6: Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=3$.
 From this table, we can draw the following conclusions: In most cases, the performance of using RDiffBR is superior to that of not using it or using RDiffBR w/o R.
 Using RDiffBR w/o R may lead to worse performance than RDiffBR, and in some cases it may even be worse than only using backbone BR model.
-Figure 6-6 show the visualization results of CrossCBR, CrossCBR+RDiffBR and CrossCBR+RDiffBR w/o R under two B-I varation levels for item-level bundle embeddings $E^{IL}_{B}$ in Youshu dataset.
+Figure 6 shows the visualization results of CrossCBR, CrossCBR+RDiffBR and CrossCBR+RDiffBR w/o R under B-I varation levels $\rho=3$ for item-level bundle embeddings $E^{IL}_{B}$ in Youshu dataset.
 It can be observed that $E^{IL}_{B}$ generated by CrossCBR after B-I change is more dispersed than original $E^{IL}_{B}$ and more compact $E^{IL}_{B}$ distribution which retains the basic shape can be obtained after using CrossCBR+RDiffBR.
@@
 While using CrossCBR+RDiffBR w/o R, $E^{IL}_{B}$ distribution changes little so the effect is limited, which is consistent with the experimental results obtained in Table 3.
-We also conducted a case study to explore the specific situations of RDiffBR and RDiffBR w/o R(see detail in section3.5).
 Overall, RDiffBR outperforms backbone models, highlighting the effectiveness of RDiffBR.
-Residual connection plays an important role in RDiffBR, and merely using a diffusion model cannot generate better item-level bundle embeddings.
-Furthermore, our task can be approached without considering the Out-Of-Distribution(OOD) which means the distribution of the training data and the test data is inconsistent.
+Residual connection plays an important role in RDiffBR, and merely using a diffusion model cannot generate better item-level bundle embeddings (The performance of $\rho=-3$ is consistent with $\rho=3$, see details in extended version).
 Although different B-I affiliations are used during model testing, the bundle theme remains, which means items within a bundle remain to follow the similar distribution before and after the change.
-This can also be verified by Figure 6-6 that the shapes of two $E^{IL}_{B}$ distributions under different B-I variation $\rho$ are similar.
-3.4. Impact of Hyperparameters
-Figure 7. The impact of loss balance factor $\lambda$
-Figure 8. The impact of residual weight $\delta$
-Figure 9. The impact of dimensionality of time step embedding $d$
-In order to achieve a more detailed understanding, we selected some hyperparameters: The loss balance factor $\lambda$, the residual weight $\delta$ and the dimensionality of time step embedding $d$.
-The results of these experiments are presented in Figure 9-9 by adjusting them within a certain range.
+This can also be verified by Figure 6 that the shapes of two $E^{IL}_{B}$ distributions under different B-I variation $\rho$ are similar.
+Therefore, with using RDiffBR, our task can be approached without considering the Out-Of-Distribution (OOD) where the distribution of the training data and the test data is inconsistent.
+In extreme cases where the distribution of $E^{IL}_{B}$ changes completely, the problem becomes equivalent to cold-start bundle recommendation (cold_start_1; cold_start_2) that the original bundle id cannot be directly associated.
+Impact of Hyperparameters
+Figure 7: The impact of loss balance factor $\lambda$ .
+Figure 8: The impact of residual weight $\delta$ .
+In order to achieve a more detailed understanding, two key hyperparameters are studied, i.e., the loss balance factor $\lambda$ and the residual weight $\delta$ .
+The results of our hyperparameter study are presented in Figure 8-8.
 $\bullet$ Loss balance factor $\lambda$.
-The loss balance factor $r$ affects the performance of RDiffBR, which is achieved by adjusting the emphasis between the recommendation task and the reconstruction task.
-A high value of $r$ may neglect the core recommendation task and thereby compromise the overall performance.
-$\bullet$ Residual weight $\delta$.
-The appropriate selection of residual weights is crucial for achieving embedding generation, this is because too high or too low values will lead to a decline in performance.
-$\bullet$ Dimensionality of time step embedding $d$.
-As the value of d increases, the model performance initially improves.
-This is because the time step embedding can flexibly adjust the denoising strategy according to different noise stages.
-However, when the value of d is excessively large, the input embedding will gradually be weakened during the calculation process, thereby negatively affecting the model performance.
-3.5. Case Study
-Figure 10. A case study for Youshu under $\rho=-3$. The highlighted items are those before the bundle change.
-Figure 11. A case study for Youshu under $\rho=3$. The highlighted items are those before the bundle change.
-In this section, we analyze the real data instances as show in Figure 11-11.
-We randomly select two bundles from Youshu and calculate cosine similarity between item-level bundle embeddings $E^{IL}_{B}$ and all items after B-I change.
+The loss balance factor $r$ affects the performance of RDiffBR, which adjusts the emphasis between the recommendation task and the reconstruction task.
+A high value of $r$ may neglect the core recommendation and thereby compromise the overall performance.
+$\bullet$ Residual weight $\delta$. The appropriate selection of residual weights is crucial for embedding generation, this is because too high or too low values will lead to a performance decline.
+Case Study
+Figure 9: A case study for Youshu under $\rho=-3$. The highlighted items are those before the bundle change.
+Figure 10: A case study for Youshu under $\rho=3$. The highlighted items are those before the bundle change.
+In this section, we analyze the real data instances as show in Figure 9.
+We randomly select a bundle from Youshu and calculate cosine similarity between item-level bundle embeddings $E^{IL}_{B}$ and all items under B-I varation levels $\rho=-3$.
 We listed the top 5 items with the highest similarity scores and their corresponding similarity values.
@@
 When using RDiffBR, the model can not only express the main items of the bundle but also introduce other items based on the theme to enhance the generalization ability.
-3.6. Training Efficiency Study (RQ3)
-In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time(seconds) of using BR model and BR model+RDiffBR on two different scale datasets.
-We record five consecutive training epochs and average them to obtain the one-epoch training time, as shown in Table 4.
+Training Efficiency Study (RQ3)
+In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes).
+We recorded five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4.
 It shows that RDiffBR only increased training time about 4%.
-Table 4. The statistics of one-epoch training time(seconds) for three backbone models and using RDiffBR on them.
-Dataset Youshu NetEase
-BGCN 1.46 23.5
-BGCN+RDiffBR 1.51 24.5
-CrossCBR 1.06 8.7
-CrossCBR+RDiffBR 1.11 9.1
-BundleGT 22.7 426.2
-BundleGT+RDiffBR 23.3 434.8
-4. RELATED WORK
-4.1. Bundle Recommendation
-Bundle recommendation aims to recommend a bundle of items related to a specific theme.
-Early work employed neural network methods for bundle recommendation (Cao et al., 2017; Chen et al., 2019b; He et al., 2019; Liu et al., 2014).
-Recently, graph neural networks have become the mainstream technology for bundle recommendation due to its strong ability to encode higher-order information.
-We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Sun et al., 2024).
-$\bullet$ Unified learning.
-The unified representation learning strategy aims to learn representations of users and bundles within a unified framework.
-There are some existing research studies that utilized this strategy, which integrated users, items and bundles into a unified user-bundle-item graph (Chen et al., 2019a) or hypergraph  (Yu et al., 2022), and propagated neighboring information to get the final representations of users and bundles.
-$\bullet$ Separate Learning.
-The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction.
-This works (Chang et al., 2020; Ke et al., 2023b; Tan et al., 2021; Li et al., 2023; Wei et al., 2023) perform representation learning on several views separately and then fuse the specific predictions of views.
-$\bullet$ Cooperative Learning.
-This approach is similar to the separate learning strategy in that it learns user and bundle representations from bundle view and item view but the key difference is that, techniques such as contrastive learning (Zhao et al., 2022; Ma et al., 2022, 2024a; Zou et al., 2024), knowledge distillation (Ke et al., 2023a; Ren et al., 2023) are employed cross views to achieve mutual enhancement.
-However, most of bundle recommendation models have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles.
-So they lack the ability to adapt to item-level dynamic variability scenario.
-It is necessary to study our task for real-world application scenarios.
-4.2. Diffusion Models for Recommendation
-The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation.
-Diffusion model plays three main roles in recommendation systems (Lin et al., 2024b): $\bullet$ Diffusion for data augmentation.
-By generating realistic samples to expand the original training data, such as sequence data (Wu et al., 2023; Liu et al., 2023), multi-field features (Kotelnikov et al., 2023; Zhang et al., 2024), user-item interactions (Liu et al., 2024) and side information (Jiang et al., 2024b, a), diffusion models can enhance the performance of recommendation systems from multiple perspectives.
-$\bullet$ Diffusion for Representation Enhancement.
-Diffusion models converts the input data into embeddings to address some key challenges of recommendation.
-Lots of studies have been proposed such as denoising the implicit user feedback (Zhao et al., 2024), out-of-distribution issue (Zhao et al., 2025) and multi-modal learning (Ma et al., 2024b).
-$\bullet$ Diffusion as Recommendation Model.
-When using diffusion models as recommendation models, the input at this stage can be structured datas or embeddings from other encoders.
-The related research can be classified into three categories: collaborative recommendation (Wang et al., 2023; Hou et al., 2024), context-aware recommendation (Li et al., 2024b; Qin et al., 2024), and other applications (Lin et al., 2024a).
-$\bullet$ Diffusion for Content Presentation.
-Diffusion models can further enhance user satisfaction by presenting generated content to target users, rather than conducting item recommendations.
-Existing research works can be roughly classified into three categories: fashion generation (Xu et al., 2024), advertising creative generation (Yang et al., 2024) and general content generation (Wang et al., 2024).
-However, simply applying diffusion model on bundle models is insufficient to address the issue of item-level dynamic variability.
-There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
-5. Conclusion and Future Work
-In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, which is a model-agnostic generative framework that can help a BR model adapt this scenario.
-Given the item-level recommendation model of the bundle, RDiffBR actively injects Gaussian noise into these embeddings and iteratively removes the noise during the reverse process.
-To guide the reverse process, we designed a residual app module.
-In the inference stage, we used this module to generate new item-level embeddings.
-A large number of experimental results show that RDiffBR enables the backbone BR model to achieve significant performance improvement in scenarios with dynamic changes of B-I.
-Future improvements of RDiffBR may include: 1) allowing RDiffBR to be directly inserted into the well-trained bundle model and take effect; 2) Using diffusion to model bundle features as distridution rather than fixed features that have limitations in B-I dynamic variability scenario.
-Acknowledgements.
+[…]
+BGCN 1.46 23.5 68.3 1.07
++RDiffBR 1.51 24.5 71.4 1.09
+Extra cost 3.42% 4.25% 4.53% 1.86%
+CrossCBR 1.06 8.70 45.4 1.10
++RDiffBR 1.11 9.10 46.2 1.14
+Extra cost 4.71% 4.59% 1.76% 3.63%
+BundleGT 22.7 426.2 118.4 1.23
++RDiffBR 23.3 434.8 121.1 1.29
+Extra cost 2.64% 2.01% 2.28% 4.87%
+Table 4: The statistics of one-epoch training time (seconds) for three backbone models and using RDiffBR on them.
+From the perspective of time complexity, taking lightweight backbone CrossCBR as an example, the time complexity of graph learning is $\mathcal{O}(|E|Kd)$, where $|E|$ is the number of all edges in U-B and U-I graphs, $K$ is the number of propagation layers, $d$ is the embedding size.
+However, the diffusion process in RDiffBR costs $\mathcal{O}(|B|d)+\mathcal{O}(|B|d)d’n$, where $|B|$ is the number of bundles, $d$’ is the embedding size of MLP in approximator, $n$ is MLP layer number.
+It shows that $|E|$ is much greater than $|B|$ (e.g., in iFashion $|E|=1679708>>|B|=27694$).
+Therefore, our RDiffBR only increases training time about $4\%$ on two GTX 4090.
+Conclusion and Future Work
+In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, a model-agnostic generative framework that can help a BR model adapt to this scenario.
+A number of experimental results show that RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I.
+Future research will consider extending RDiffBR as a plug-and-play module for well-performing pre-trained bundle recommenders and evaluating its applicability to cold-start bundle recommendation.
+Acknowledgments
+This work is partially supported by NSFC, China (No.62276196), the research grant (RGPIN-2020-07157) from the Natural Science and Engineering Research Council (NSERC) of Canada and York Research Chairs (YRC) program.
+The authors also gratefully appreciate the anonymous reviewers for their valuable comments and constructive suggestions that greatly helped to improve the quality of the paper.

### v4

@@
 Our empirical study demonstrates that the performance of mainstream BR models may fluctuate or decline under item-level variability.
-This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario.
+This paper makes the first attempt to address the above problem and proposes Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario.
 During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by the BR model to represent bundle theme via a forward-reverse process.
@@
 In particular, the residual connection in our residual approximator significantly enhances BR models’ ability to generate high-quality item-level bundle embeddings.
-Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
+Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increases training time about 4%.
 Codehttps://github.com/WUT-IDEA/RDiffBR
@@
 For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases (see Section 3.2 for details).
-This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations regarding item-level.
+This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations of item-level.
 In this paper, we make the first attempt to address the issue of item-level dynamic variability in bundle recommendation and propose a novel Residual Diffusion for Bundle Recommendation (RDiffBR) model-agnostic generative framework to assist the model in adapting this scenario.
@@
 This approach learns user and bundle representations from bundle view and item view and uses techniques such as contrastive learning (MIDGN; CrossCBR; MultiCBR; HIDGN), knowledge distillation (HyperMBR; DGMAE) or diffusion (DCBR) to achieve cross views mutual enhancement.
-However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability scenario.
+However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability.
 It is necessary to study our task for real-world application scenarios.
@@
 However, a straightforward application of diffusion models on bundle recommendation is insufficient to address the issue of item-level dynamic variability.
-There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
+There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistent and the approximator needs to be designed based on item-level dynamic variability scenario.
 Methodology
@@
 Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation.
-And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
+When an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
 RDiffBR Framework
@@
 […]
-$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
+$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarly factorize $p_{\theta}(e_{t-1}|e_{t})$:
 $p_{\theta}(\hat{e}_{t-1}|\hat{e}_{t})=\mathcal{N}(\hat{e}_{t-1};\mu_{\theta}(\hat{e}_{t},t),\tilde{\beta}_{t}\boldsymbol{I})),\quad\mathrm{~where}$ (6)
@@
 Training Efficiency Study (RQ3)
-In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes).
-We recorded five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4.
+In order to evaluate the additional computing time of RDiffBR, we compare the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes).
+We record five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4.
 It shows that RDiffBR only increased training time about 4%.
@@
 Conclusion and Future Work
-In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, a model-agnostic generative framework that can help a BR model adapt to this scenario.
-A number of experimental results show that RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I.
+In this work, to address the issue of item-level dynamic variability, we propose RDiffBR, a model-agnostic generative framework that can help BR models adapt to this scenario.
+A number of experiments show RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I.
 Future research will consider extending RDiffBR as a plug-and-play module for well-performing pre-trained bundle recommenders and evaluating its applicability to cold-start bundle recommendation.